### v2

# Fine-grained Image Quality Assessment for Perceptual Image Restoration

Abstract
Recent years have witnessed remarkable achievements in perceptual image restoration (IR), creating an urgent demand for accurate image quality assessment (IQA), which is essential for both performance comparison and algorithm optimization. Unfortunately, the existing IQA metrics exhibit inherent weakness for IR task, particularly when distinguishing fine-grained quality differences among restored images. To address this dilemma, we contribute the first-of-its-kind fine-grained image quality assessment dataset for image restoration, termed FGRestore, comprising 18,408 restored images across six common IR tasks. Beyond conventional scalar quality scores, FGRestore was also annotated with 30,886 fine-grained pairwise preferences. Based on FGRestore, a comprehensive benchmark was conducted on the existing IQA metrics, which reveal significant inconsistencies between score-based IQA evaluations and the fine-grained restoration quality. Motivated by these findings, we further propose FGResQ, a new IQA model specifically designed for image restoration, which features both coarse-grained score regression and fine-grained quality ranking. Extensive experiments and comparisons demonstrate that FGResQ significantly outperforms state-of-the-art IQA metrics. Codes and model weights have been released in https://pxf0429.github.io/FGResQ/.
Introduction
Perceptual image restoration (IR) stands as a cornerstone in low-level computer vision, aiming to recover high-quality images from their degraded observations while maintaining perceptual fidelity (Potlapalli et al. 2023; Luo et al. 2023). Recent years have witnessed remarkable breakthroughs in this field, largely driven by the evolution of generative models. These advances have enabled IR algorithms to achieve unprecedented visual quality, creating an urgent demand for accurate image quality assessment (IQA) methods that can reliably evaluate and compare restored images. Such assessment capabilities are essential not only for performance benchmarking across different IR algorithms but also for guiding algorithmic optimization.
Currently, IR algorithm evaluation still predominantly relies on traditional reference-based metrics such as PSNR and SSIM (Wang et al. 2004), which assess restoration quality by measuring the similarity between the restored images and ground-truth images, even though their inconsistency with human perceptual judgment has been increasingly observed by the research community (Jinjin et al. 2020). Recent efforts have introduced no-reference IQA methods (Xu et al. 2024; Zhou et al. 2025), including CLIP-IQA (Wang, Chan, and Loy 2023) and other learning-based approaches (Ke et al. 2021; Wu et al. 2023), which prove particularly valuable in real-world restoration scenarios where ground truth images are unavailable. However, a critical oversight persists: these methods have been primarily designed for image quality assessment, without specific consideration for the fine-grained evaluation requirements inherent in IR tasks.
Figure 1: Illustration of the fine-grained challenge in IQA for image restoration. Both IR algorithm comparison and optimization processes require distinguishing subtle quality differences between restored images. Existing IQA metrics fail to provide correct rankings for fine-grained image pairs. (Best viewed zoomed in.)
Figure 2: Overview of our method. (a) FGRestore provides comprehensive fine-grained quality annotations for multiple IR tasks. (b) FGResQ enables both coarse-grained quality scoring and fine-grained quality ranking capabilities.
A critical yet underexplored challenge in IR evaluation lies in its inherently fine-grained nature. As illustrated in Figure 1, IR algorithm comparison and optimization typically involve distinguishing between images with subtle quality differences—scenarios where even state-of-the-art IQA methods struggle to provide reliable assessments. This observation leads us to fundamentally rethink existing evaluation paradigms for IR tasks and raises a crucial question:
Can existing score-based IQA methods objectively capture fine-grained quality differences in restored images?
To investigate this question, we first conducted comprehensive computational analyses on established IQA datasets for image restoration (Jinjin et al. 2020; Shi et al. 2019). Our findings consistently reveal a significant performance gap: while existing IQA methods achieve reasonable results in coarse-grained quality assessment, they consistently underperform in fine-grained scenarios, with significant inconsistencies between predicted scores and human perceptual judgments. This observation highlights the inadequacy of current evaluation protocols for capturing the nuanced quality differences which are crucial in IR applications. Moreover, when quality differences between images become subtle, both human and IQA models encounter significant challenges in providing reliable absolute quality scores. In contrast, pairwise comparison presents a more reliable alternative for fine-grained assessment, as humans demonstrate superior consistency in relative quality judgments compared to absolute scoring.
Motivated by the above insights, we introduce FGRestore, the first fine-grained IQA dataset specifically for image restoration tasks, as shown in Figure 2(a). Our dataset comprises 18,408 restored images spanning six common IR tasks: image super-resolution, deblurring, denoising, dehazing, deraining, and mixture restoration. Beyond conventional Mean Opinion Score (MOS) annotations, FGRestore incorporates 30,886 fine-grained preference annotations for restored image pairs with subtle quality differences. Based on FGRestore, we further propose FGResQ, a unified fine-grained IQA model tailored for IR evaluation (Figure 2(b)). Our approach employs degradation-aware feature learning that incorporates human-level degradation knowledge from Vision-Language Models (VLMs) into the proposed IQA framework, enabling both accurate quality scoring and fine-grained quality ranking across diverse IR tasks.
Our main contributions are threefold:
  • We construct FGRestore, the first fine-grained IQA dataset specifically designed for IR tasks, containing 18,408 restored images across six common IR categories with both MOS and 30,886 fine-grained pairwise preference annotations.
  • We propose FGResQ, a unified fine-grained IQA model that incorporates degradation-aware feature learning and demonstrates state-of-the-art performance in both coarse-grained score regression and fine-grained ranking.
  • Through comprehensive computational analysis on existing IR datasets and extensive experiments, we reveal fundamental limitations of current score-based IQA methods in capturing fine-grained quality differences, providing valuable insights for future IR evaluation research.
Related Work
Image Restoration Quality Assessment Dataset
To evaluate IR algorithms, numerous datasets have been proposed. These include datasets dedicated to single restoration types, such as the MDD13 (Liu et al. 2013) dataset for deblurring, the IVC-Dehazing (Ma, Liu, and Wang 2015) and exBeDDE (Zhao et al. 2020) datasets for dehazing, the IVIPC-DQA (Wu et al. 2019) dataset for deraining, and the QADS (Zhou et al. 2019), SISRSet (Shi et al. 2019), RealSRQ (Jiang et al. 2022) and SRIQA-Bench (Chen et al. 2025) datasets for super-resolution. Additionally, there are datasets that encompass distortions from various algorithms, such as PIPAL (Jinjin et al. 2020), which includes distortion types like denoising, super-resolution, and super-resolution with post-denoising, KADID-10k (Lin, Hosu, and Saupe 2019), which contains various image distortions including denoising degradations and DiffIQA(Chen et al. 2025), which accounts for degradation artifacts induced by diffusion-based image enhancement. A brief comparison of FGRestore and previous IR evaluation datasets is illustrated in Table 1.
Image Quality Assessment
Existing image restoration (IR) algorithms typically employ full-reference image quality assessment (FR-IQA) metrics for performance evaluation, such as the conventional PSNR and SSIM (Wang et al. 2004), as well as deep learning-based metrics like LPIPS (Zhang et al. 2018a) and DISTS (Ding et al. 2020). To overcome the flawed ”perfect reference” assumption in traditional FR-IQA, which is challenged by imaging system limitations and superior generative methods, A-FINE (Chen et al. 2025) proposes a generalized model that adaptively assesses both the fidelity and naturalness of test images.
No-reference IQA (NR-IQA) methods have been proposed to assess image quality without reference images. Early NR-IQA methods, such as NIQE (Mittal, Soundararajan, and Bovik 2012), primarily rely on hand-crafted Natural Scene Statistics (NSS) features. With the advancement of deep learning, many works have turned to neural networks for NR-IQA. For instance, MetaIQA (Zhu et al. 2020) leverages meta-learning to enhance model generalization. Additionally, methods such as CLIP-IQA (Wang, Chan, and Loy 2023) leverage text prompts to enhance perceptual capabilities. Given the powerful learning capabilities of large multi-modality models (LMMs), emerging work, such as Q-Align (Wu et al. 2023), has begun to apply these models to quality assessment. Most recently, methods represented by Co-Instruct (Wu et al. 2024) have extended LMMs to compare multiple images. Furthermore, VisualQuality-R1(Wu et al. 2025) introduces a novel methodology for IQA that leverages LMMs within a reinforcement learning-to-rank framework.
Although these datasets and methods have contributed significantly to the IQA community, most methods are score-based and cannot effectively handle fine-grained perceptual differences brought by image restoration algorithms. Recent works have explored fine-grained quality assessment in specific domains. Zhang et al. (Zhang et al. 2018b; Zhang, Lin, and Huang 2021) introduced fine-grained assessment for compressed images, focusing on distinguishing compression artifacts within similar quality levels. However, these prior works primarily target image compression, whereas our work specifically addresses the unique challenges of fine-grained quality assessment in image restoration tasks, where the evaluation focus shifts from compression artifacts to perceptual restoration quality differences.
Table 1: Comparison with the previous datasets. MOS: Mean Opinion Score; PC: pairwise comparison; ACR: Absolute Category Rating; DCR: Degradation Category Rating; 5-level: 5-point quality scale. Number show Images/Pairs/Annotations.
Dataset Year Task Alg. Annotation Number
Type (Img./Pairs/Ann.)
MDD13 2013 1 5 MOS (PC) 1,200/0/13,592
IVC-Dehazing 2015 1 8 MOS (ACR) 200/0/4,800
QADS 2019 1 21 MOS (PC) 1,260/0/126,000
SISRSet 2019 1 8 MOS (PC) 260/0/32,000
IVIPC-DQA 2019 1 6 MOS (5-level) 2,136/0/27,192
KADID-10k 2019 1 1 MOS(DCR) 10,125/0/303,750
exBeDDE 2020 1 10 MOS (PC) 1,670/0/18,380
PIPAL 2020 3 40 MOS (Elo) 29,000/0/1.13M
RealSRQ 2022 1 10 MOS (PC) 1,620/0/65,400
DiffIQA 2025 1 1 RANK 177,319/177,319/537,624
SRIQA-Bench 2025 1 10 RANK 1100/5500/55000
FGRestore 2025 6 108 MOS+RANK 18,408/30,886/45,318
Table 2: Performance comparison across different MOS ranges on PIPAL dataset.
Type Method [0.0,0.2) [0.2,0.4) [0.4,0.6) [0.6,0.8) [0.8,1.0] Overall
SRCC SRCC SRCC SRCC SRCC SRCC
FR PSNR 0.323 0.082 0.209 0.161 0.072 0.422
SSIM 0.293 0.108 0.258 0.254 0.049 0.530
LPIPS -0.034 0.077 0.325 0.287 0.124 0.612
DISTS 0.168 0.159 0.310 0.242 0.165 0.585
NR NIQE -0.126 -0.002 0.107 0.001 0.080 0.153
IL-NIQE -0.235 -0.098 0.126 0.128 0.054 0.289
BRISQUE -0.142 0.025 0.125 0.035 0.131 0.185
DB-CNN -0.157 0.321 0.353 0.330 -0.016 0.636
HyperIQA 0.100 0.274 0.314 0.292 0.032 0.584
MetaIQA 0.037 0.160 0.204 0.174 -0.101 0.423
LIQE -0.232 0.053 0.175 0.299 0.107 0.479
CLIP-IQA -0.152 0.211 0.238 0.293 0.071 0.530
Q-Align 0.230 0.301 0.337 0.213 0.178 0.418
DeQA-Score 0.568 0.676 0.623 0.516 0.350 0.747
Figure 3: Overview of the FGRestore dataset construction.
Preliminary Validation Analysis
In image restoration tasks, both algorithm comparison and optimization frequently involve evaluating images with subtle quality differences. Algorithm comparison requires distinguishing between restoration results with marginal quality differences, while parameter optimization involves incremental quality changes that demand sensitive assessment methods to identify optimal configurations. To investigate whether existing IQA methods can objectively capture fine-grained quality differences in IR task, we conduct a comprehensive computational analysis on established IQA datasets. Specifically, we evaluate state-of-the-art IQA methods to assess their fine-grained discrimination capabilities. Our evaluation encompasses both full-reference (FR) and no-reference (NR) methods. FR methods include PSNR, SSIM (Wang et al. 2004), LPIPS (Zhang et al. 2018a), and DISTS (Ding et al. 2020). NR methods include traditional approaches such as NIQE (Mittal, Soundararajan, and Bovik 2012), IL-NIQE (Zhang, Zhang, and Bovik 2015), and BRISQUE (Mittal, Moorthy, and Bovik 2012), as well as deep learning-based methods including DB-CNN (Zhang et al. 2020), HyperIQA (Su et al. 2020), MetaIQA (Zhu et al. 2020), LIQE (Zhang et al. 2023), CLIP-IQA (Wang, Chan, and Loy 2023), Q-Align (Wu et al. 2023), and DeQA-Score (You et al. 2025).
Specifically, we partition the quality score range into several intervals to examine how IQA methods perform when evaluating images with similar quality levels, which is particularly relevant for fine-grained assessment scenarios in image restoration. Table 2 presents the performance comparison across different MOS ranges on the PIPAL dataset (Jinjin et al. 2020). While most IQA methods achieve reasonable overall SRCC values on the whole dataset, their performance dramatically deteriorates when evaluated within narrow quality ranges. We also conducted similar analyses on other IQA datasets, which consistently demonstrate the same pattern. More results are provided in the supplementary material. These findings provide strong evidence for the inadequacy of current score-based IQA approaches in fine-grained scenarios, motivating the development of our fine-grained evaluation framework for image restoration.
FGRestore Dataset
Motivated by the findings revealed in our preliminary analysis, we construct FGRestore, the first comprehensive fine-grained image quality assessment dataset for perceptual image restoration evaluation. The dataset construction pipeline is illustrated in Figure 3, which involves collecting images from multiple datasets, filtering image pairs through score-based and JND-based criteria to retain fine-grained pairs, and conducting systematic subjective studies for pairwise preference annotation.
Image and Pair Collection
We collect images from multiple restoration-specific datasets to cover different degradation and visual appearances: 1) deraining images from IVIPC-DQA (Wu et al. 2019); 2) deblurring images from MDD13 (Liu et al. 2013); 3) dehazing images from IVC-Dehazing (Ma, Liu, and Wang 2015) and exBeDDE (Zhao et al. 2020); 4) denoising images from PIPAL (Jinjin et al. 2020) and KADID-10k (Lin, Hosu, and Saupe 2019); 5) super-resolution images from QADS (Zhou et al. 2019), RealSRQ (Jiang et al. 2022), SISRSet (Shi et al. 2019), and PIPAL (Jinjin et al. 2020); 6) mixture restoration images from PIPAL (Jinjin et al. 2020). After image collection, we generate image pairs by randomly pairing all images with identical content and restoration tasks, resulting in a total of 1,275,297 image pairs.
Data Filtration
The randomly generated image pairs can be categorized into three distinct types based on perceptual quality differences: (1) Coarse-grained pairs with highly noticeable quality differences, (2) Unnoticeable pairs with negligible quality differences, and (3) Fine-grained pairs with subtle but perceptible quality differences. We employ a two-step filtration process to remove coarse-grained and unnoticeable pairs.
Score-based Coarse-grained Pairs Filtering.
To eliminate coarse-grained pairs, we establish score difference thresholds for each source dataset through expert discussion. The coarse-grained filtering criterion is defined as:
$\mathcal{F}_{\text{coarse}}(p_{i})=\begin{cases}1,&\text{if }|s_{i}^{A}-s_{i}^{B}|\leq\tau_{d}\\ 0,&\text{otherwise}\end{cases},$ (1)
where $p_{i}=(I_{i}^{A},I_{i}^{B})$ represents the $i$-th image pair, $s_{i}^{A}$ and $s_{i}^{B}$ are the quality scores of images $I_{i}^{A}$ and $I_{i}^{B}$ respectively, and $\tau_{d}$ is the dataset-specific threshold. Pairs with $\mathcal{F}_{\text{coarse}}(p_{i})=1$ are retained for further processing.
JND-based Unnoticeable Pairs Filtering.
To remove unnoticeable pairs, we propose a JND-based filtering approach. The JND represents the minimum perceptual threshold for noticeable quality differences, making it ideal for identifying imperceptible quality pairs. Our JND-based filtering process operates as follows: (1) randomly select 20,000 images from the collected datasets, (2) compute the JND map for each image using established JND estimation model (Wu et al. 2017), and (3) generate JND noise images by adding these JND maps as noise to the original images:
$I_{\text{JND}}=I+\text{JND}(I),$ (2)
where $I$ is the original image and $\text{JND}(I)$ is the corresponding JND noise map. We calculate SSIM (Wang et al. 2004) between each $I$ and $I_{\text{JND}}$, and use the median SSIM value $\text{SSIM}_{\text{med}}$ as our filtering threshold. The unnoticeable pair filtering criterion is then defined as:
$\mathcal{F}_{\text{unnotice}}(p_{j})=\begin{cases}1,&\text{if }\text{SSIM}(I_{j}^{A},I_{j}^{B})\leq\text{SSIM}_{\text{med}}\\ 0,&\text{otherwise}\end{cases},$ (3)
where $p_{j}=(I_{j}^{A},I_{j}^{B})$ represents an image pair that passed the coarse-grained filtering. Pairs with $\mathcal{F}_{\text{unnotice}}(p_{j})=1$ are retained as fine-grained pairs.
After applying both filtering steps, we retain 30,886 fine-grained image pairs and their corresponding 18,408 images, forming the core of our FGRestore dataset. Due to space limitations in the main paper, we are unable to clearly demonstrate fine-grained restoration image pair examples at sufficient resolution. Additional examples of fine-grained restoration image pairs are provided in the supplementary material.
In-lab Subjective Study
FGRestore preserves the original score annotations from source datasets while introducing comprehensive fine-grained pairwise ranking annotations. For datasets originally annotated through pairwise ranking methodologies, fine-grained rankings can be directly derived from existing score relationships without additional human annotation. We design a resource-efficient two-round annotation protocol that incorporates quality control mechanisms to ensure reliable annotations. The annotation process is structured as follows:
Annotation Protocol. (1) Image pairs are divided into two groups, with each group assigned to a team of five trained annotators. For each image pair, annotators select the image with superior perceptual quality. Recognizing that some pairs may contain imperceptible differences despite filtering, annotators can also indicate equal quality when no clear preference exists. (2) Due to the subjective characteristics of fine-grained quality assessment, disagreements inevitably arise between annotators when quality differences are particularly subtle. Inconsistent annotations are resolved through expert review. This systematic annotation approach yields 45,318 pairwise preference annotations, providing a robust foundation for fine-grained quality assessment model training and evaluation.
Dataset Analysis
To validate the necessity of fine-grained pairwise ranking annotations, we conduct a consistency analysis between MOS-based scoring and human preference ranking. We create a scatter plot comparing normalized MOS differences ($x$-axis) with human preference annotations ($y$-axis). As shown in Figure 4, the scatter plot reveals significant inconsistencies between MOS scores and human preferences. Points in the red regions represent inconsistent annotations where MOS scores and ranking preferences disagree. Critically, image pairs with smaller MOS differences exhibit larger scatter points, indicating that inconsistencies become more severe as quality differences become subtler. This strongly supports the necessity of fine-grained pairwise ranking, as MOS-based scoring becomes unreliable for discriminating subtle quality differences.
Figure 4: Consistency analysis between MOS scores and human preference rankings. Point sizes represent frequency of image pairs. Red regions indicate inconsistent cases where MOS scores and human preferences disagree.
Figure 5: Overview of the proposed FGResQ framework.
FGResQ
Based on FGRestore, we propose FGResQ, a new fine-grained image quality assessment model for perceptual image restoration evaluation. The overall framework pipeline is illustrated in Figure 5, which consists of two main components: (a) Degradation-aware Feature Learning that incorporates restoration task knowledge to enable unified evaluation across multiple IR tasks, and (b) Dual-branch Quality Prediction that simultaneously handles both coarse-grained score regression and fine-grained pairwise ranking.
Degradation-aware Feature Learning
To achieve unified quality assessment across diverse IR tasks, we propose a degradation-aware feature learning approach that enables our model to perceive and incorporate task-specific degradation characteristics. This design allows FGResQ to handle multiple IR scenarios within a unified framework, eliminating the need for task-specific model training. Specifically, we leverage a pretrained CLIP model to establish semantic alignment between visual content and degradation type. Specifically, we freeze the text encoder while fine-tuning a degradation encoder to learn degradation-aware representations. The degradation encoder maps images to a feature space that distinguishes different restoration scenarios. The learning objective employs bidirectional contrastive alignment to ensure robust degradation type perception:
$\mathcal{L}_{cont}=\frac{1}{2}\left[\mathcal{L}_{CE}(\mathbf{F}_{I}\mathbf{F}_{T}^{T},\mathbf{y})+\mathcal{L}_{CE}(\mathbf{F}_{T}\mathbf{F}_{I}^{T},\mathbf{y})\right],$ (4)
where $\mathbf{F}_{I}$ and $\mathbf{F}_{T}$ represent image and text features respectively, $\mathcal{L}_{CE}$ denotes the cross-entropy loss, and $\mathbf{y}=[0,1,2,\ldots,N-1]$ represents the ground-truth matching labels for $N$ samples in a batch.
Dual-branch Quality Prediction
Given degradation-aware features from the previous module, we first enhance them with learnable prompt embeddings to obtain quality-aware representations, then employ specialized prediction heads for different assessment tasks. Given an input image $I$, we extract general features $\mathbf{F}_{g}\in\mathbb{R}^{d}$ using the image encoder and degradation features $\mathbf{F}_{d}\in\mathbb{R}^{d}$ from the frozen degradation encoder. To effectively utilize degradation information, we employ learnable prompt embeddings $\mathbf{p}\in\mathbb{R}^{d}$ to transform degradation features into quality-aware features:
$\begin{aligned} \mathbf{F}_{p}&=\text{MLP}_{1}(\text{softmax}(\text{MLP}_{2}(\mathbf{F}_{d}))\cdot\mathbf{p})\\ \mathbf{F}_{q}&=\mathbf{F}_{g}\oplus\mathbf{F}_{p}\oplus(\mathbf{F}_{g}+\mathbf{F}_{p})\\ \end{aligned},$ (5)
where $\text{MLP}_{1}$ and $\text{MLP}_{2}$ are multi-layer perceptrons, $\oplus$ denotes concatenation, and $\mathbf{F}_{q}$ represents the final quality-aware features.
FGResQ employs two specialized prediction heads: Regression Head processes the quality features $\mathbf{F}_{q}$ to predict absolute quality scores $y_{pred}$ for individual images, enabling coarse-grained quality assessment compatible with traditional evaluation protocols. Comparison Head handles fine-grained quality ranking by processing image pairs $(I_{A},I_{B})$ to obtain their respective quality features $(\mathbf{F}_{q}^{A},\mathbf{F}_{q}^{B})$ and predicting the ranking probability $p_{AB}$, directly addressing the fine-grained discrimination challenges identified in our preliminary analysis.
Training Objectives
We design a comprehensive training strategy through carefully designed loss functions:
Scene-aware Fidelity Loss. Since FGRestore preserves original score annotations from different source datasets with varying quality scales, we employ a scene-aware fidelity loss that focuses on ranking relationships within each dataset (scene) rather than absolute score values:
$\mathcal{L}_{fid}^{(s)}$ $=\frac{1}{N_{s}(N_{s}-1)/2}\sum_{i<j}\Big{(}1-\sqrt{p_{ij}g_{ij}+\epsilon}$
$\quad-\sqrt{(1-p_{ij})(1-g_{ij})+\epsilon}\Big{)},$ (6)
where $p_{ij}=\sigma(y_{pred,i}-y_{pred,j})$, $g_{ij}=\frac{1}{2}(\text{sign}(y_{gt,i}-y_{gt,j})+1)$, and $N_{s}$ is the number of samples in scene $s$.
The overall scene-aware loss combines fidelity losses from different scenes with inverse sample count weighting:
$\mathcal{L}_{scene}=\sum_{s\in\mathcal{S}}w_{s}\mathcal{L}_{fid}^{(s)},$ (7)
where $w_{s}$ is the inverse sample count weight for scene $s$. Pairwise Ranking Loss. For fine-grained pairwise preferences, we employ binary cross-entropy loss:
$\mathcal{L}_{rank}=-\frac{1}{M}\sum_{i=1}^{M}[r_{i}\log(p_{i})+(1-r_{i})\log(1-p_{i})],$ (8)
where $M$ is the number of image pairs, $r_{i}$ is the ground-truth ranking label, and $p_{i}$ is the predicted ranking probability. The overall training objective combines both components:
$\mathcal{L}_{total}=\lambda_{1}\mathcal{L}_{scene}+\lambda_{2}\mathcal{L}_{rank},$ (9)
where $\lambda_{1}$ and $\lambda_{2}$ are balancing hyperparameters.
Table 3: Performance comparison on FGRestore dataset across different IR tasks. ”-” indicates no reference images available.
Type Method Pub. Deblurring Denoising Deraining
SRCC PLCC ACC SRCC PLCC ACC SRCC PLCC ACC
FR PSNR - 0.187 0.167 0.634 0.487 0.482 0.775 - - -
SSIM TIP’04 0.441 0.348 0.695 0.642 0.652 0.789 - - -
LPIPS CVPR’18 0.776 0.700 0.755 0.673 0.680 0.765 - - -
DISTS TPAMI’20 0.907 0.901 0.845 0.679 0.672 0.739 - - -
A-FINE CVPR’25 0.907 0.796 0.865 0.624 0.620 0.747 - - -
NR NIQE SPL’12 0.382 0.410 0.529 0.240 0.151 0.564 0.030 0.057 0.694
BRISQUE TIP’12 0.354 0.354 0.555 0.254 0.114 0.569 0.112 0.102 0.603
DB-CNN TCSVT’22 0.788 0.786 0.688 0.478 0.431 0.611 0.243 0.259 0.437
HyperIQA CVPR’25 0.871 0.887 0.402 0.605 0.625 0.675 0.264 0.294 0.484
CLIP-IQA AAAI’23 0.867 0.785 0.765 0.474 0.440 0.625 0.241 0.221 0.349
Q-Align ICML’24 0.767 0.804 0.795 0.676 0.687 0.731 0.433 0.421 0.455
DeQA-Score CVPR’25 0.815 0.843 0.819 0.754 0.771 0.778 0.507 0.576 0.426
Compare2Score NeurIPS’24 0.769 0.813 0.757 0.661 0.679 0.687 0.074 0.108 0.790
FGResQ - 0.926 0.910 0.873 0.759 0.777 0.760 0.496 0.518 0.778
Method Dehazing MixtureRestoration SuperResolution Average
SRCC PLCC ACC SRCC PLCC ACC SRCC PLCC ACC SRCC PLCC ACC
PSNR - - - 0.280 0.248 0.643 0.296 0.303 0.624 0.313 0.300 0.669
SSIM - - - 0.421 0.379 0.684 0.351 0.361 0.641 0.464 0.435 0.702
LPIPS - - - 0.466 0.475 0.658 0.448 0.460 0.666 0.591 0.579 0.711
DISTS - - - 0.495 0.464 0.644 0.482 0.488 0.658 0.640 0.631 0.721
A-FINE - - - 0.536 0.528 0.645 0.421 0.434 0.688 0.622 0.594 0.736
NIQE 0.036 0.063 0.446 0.041 0.027 0.541 0.176 0.072 0.503 0.151 0.130 0.546
BRISQUE 0.132 0.088 0.436 0.201 0.064 0.575 0.142 0.040 0.514 0.199 0.127 0.542
DB-CNN 0.643 0.645 0.524 0.415 0.460 0.614 0.459 0.454 0.643 0.504 0.506 0.586
HyperIQA 0.643 0.674 0.409 0.523 0.535 0.499 0.437 0.433 0.538 0.557 0.574 0.501
CLIP-IQA 0.547 0.499 0.546 0.302 0.300 0.580 0.244 0.186 0.571 0.446 0.405 0.573
Q-Align 0.715 0.765 0.584 0.569 0.571 0.658 0.376 0.366 0.662 0.589 0.603 0.648
DeQA-Score 0.762 0.803 0.644 0.669 0.679 0.697 0.561 0.573 0.718 0.678 0.707 0.680
Compare2Score 0.334 0.381 0.436 0.494 0.519 0.635 0.317 0.315 0.607 0.441 0.469 0.652
FGResQ 0.821 0.854 0.669 0.698 0.706 0.713 0.521 0.536 0.721 0.703 0.717 0.752
Experiments
Evaluation Protocol
We compare the proposed FGResQ model against state-of-the-art FR and NR IQA methods. For fair comparison, all learning-based methods with public source codes are retrained using the same training-testing protocol with our scene-aware loss replacing their original regression loss, except for Compare2Score (Zhu et al. 2024) which operates in a zero-shot manner without training. We evaluate performance using SRCC and PLCC for quality score prediction, and accuracy (ACC) for fine-grained pairwise ranking. We split the dataset at an 8:2 ratio by image pairs. Individual image division following the corresponding pair.
Implementation Details
We adopt CLIP ViT-B/16 as the backbone encoder for both degradation-aware feature learning and dual-branch quality prediction. Training images are resized to $256\times 256$ pixels and then randomly cropped to $224\times 224$ patches. The model is trained using the Adam optimizer with cosine annealing scheduling, where the maximum learning rate is set to $5\times 10^{-6}$, batch size to 64. To ensure training stability, we employ scene-aware sampling strategy that groups samples from the same source dataset within each batch. $\lambda_{1}$ and $\lambda_{2}$ are set to 5 and 1, respectively. The model is trained for 6 epochs using all image pairs from the training set on NVIDIA RTX4090 GPUs.
Performance Evaluation
Overall Performance Analysis. Table 3 presents comprehensive performance comparisons across all six IR tasks in our FGRestore dataset. FGResQ achieves state-of-the-art performance across most evaluation metrics, with our method demonstrating particularly significant advantages in fine-grained ranking accuracy (ACC), which serves as the most critical metric for fine-grained quality assessment. The substantial improvement in ACC demonstrates FGResQ’s superior capability in distinguishing subtle quality differences. When compared to the strongest baseline methods, FGResQ shows consistent improvements across different evaluation paradigms. Among existing approaches, DeQA-Score exhibits the best coarse-grained regression performance, yet FGResQ surpasses it by significant margins with 2.5% SRCC improvement and 7.2% ACC improvement. More remarkably, our proposed method outperforms FR methods like DISTS despite these methods having access to additional reference images for quality assessment. This superior performance highlights the critical importance of explicitly modeling fine-grained quality relationships rather than relying solely on reference-based similarity measures, demonstrating that specialized fine-grained assessment capabilities are crucial for accurate IR evaluation.
Fine-grained vs. Coarse-grained Assessment. A critical observation is the performance gap between regression metrics and ranking accuracy across different methods. While some methods achieve reasonable regression performance, their fine-grained ranking capabilities remain limited.
Figure 6: Qualitative comparison. (Best viewed zoomed in.)
Qualitative Analysis. Figure 6 presents qualitative comparisons on representative fine-grained image pairs across different restoration tasks. The results demonstrate systematic limitations of existing IQA methods in fine-grained scenarios. Traditional metrics like PSNR and SSIM often produce nearly identical scores, failing to provide meaningful quality discrimination. Advanced learning-based methods such as CLIP-IQA and DeQA-Score also struggle with subtle quality differences, frequently producing incorrect rankings. In contrast, FGResQ consistently identifies the superior image across tested cases, demonstrating the effectiveness of our pairwise comparison approach. Additional qualitative examples are provided in the supplementary material.
Conclusion
In this work, we have addressed a critical yet underexplored challenge in image restoration evaluation: the inadequacy of existing IQA methods for fine-grained quality assessment. To address this fundamental limitation, we introduced FGRestore, the first fine-grained IQA dataset specifically designed for IR evaluation, comprising 18,408 restored images across six restoration tasks with 30,886 fine-grained pairwise preference annotations. Based on this, our proposed FGResQ model achieves state-of-the-art with significant improvements in fine-grained ranking accuracy. This work establishes a new evaluation framework for image restoration and provides valuable insights for developing more perceptually-aligned quality assessment methods in the era of advanced generative restoration models.
Supplementary Material for Fine-grained Image Quality Assessment for Perceptual Image Restoration
This supplementary document provides comprehensive experimental evidence, detailed analyses, and extensive visual demonstrations that substantiate the findings presented in our main paper. The material herein offers deeper insights into the fine-grained challenges of image restoration quality assessment and validates the effectiveness of our proposed approach across diverse scenarios.
This document encompasses the following components:
• Detailed Ablation Study Results: Systematic investigation of our degradation-aware feature learning (DFL) module’s contribution, demonstrating consistent improvements across multiple evaluation metrics (SRCC, PLCC, ACC) and providing empirical justification for our architectural design choices.
• Comprehensive Performance Analysis Across Quality Ranges: Exhaustive evaluation of state-of-the-art methods across 13 datasets, revealing consistent performance degradation of existing approaches within narrow MOS score intervals and highlighting the fundamental challenges in fine-grained quality assessment.
• Exemplars of Fine-Grained Quality Differences: Carefully curated image pairs demonstrating subtle quality variations across six restoration tasks - deblurring, denoising, deraining, super-resolution, dehazing, and mixture restoration. These examples illustrate the perceptual challenges that make fine-grained assessment particularly demanding for image quality assessment.
• Thorough Qualitative Analysis: Side-by-side visual comparisons showcasing how different quality assessment methods respond to restoration results with nuanced quality differences. These comparisons reveal the limitations of traditional metrics and the inconsistent behavior of deep learning-based approaches in fine-grained evaluation scenarios, while demonstrating our proposed FGResQ method in accurately identifying quality preferences that align with human perceptual judgments.
All source code and implementation details are provided in the supplementary material.
Appendix A Ablation Study
As shown in Table 3 in main paper, , unlike task-specific methods that excel in particular scenarios, FGResQ maintains robust performance across all six IR tasks, suggesting that incorporating degradation knowledge enables more generalizable quality assessment capabilities.To further validate the effectiveness of our degradation-aware feature learning (DFL) module, we conduct ablation experiments. Table 4 presents the average performance across all IR tasks, comparing our full FGResQ model with a variant that removes the degradation-aware learning component (w/o DFL).
Table 4: Ablation study.
Method SRCC PLCC ACC
w/o DFL 0.698 0.711 0.743
FGResQ 0.703 0.717 0.752
Appendix B Comprehensive Performance Analysis Across Quality Ranges
Following our methodology in the main paper, we partition quality score ranges into discrete intervals, and systematically evaluate existing IQA methods across 13 benchmark datasets. This comprehensive analysis encompasses both full-reference (FR) methods including PSNR, SSIM, LPIPS, and DISTS, as well as no-reference (NR) methods ranging from traditional approaches (NIQE, IL-NIQE, BRISQUE) to state-of-the-art learning-based methods (DB-CNN, HyperIQA, MetaIQA, LIQE, CLIP-IQA, Q-Align, DeQA-Score). Tables 2 through 14 present comprehensive results across two categories of datasets. General IQA datasets include TID2013, CSIQ, LIVE, KADID-10k and VCL@FER for synthetic distortions. Image restoration quality assessment datasets encompass PIPAL for super-resolution, denoise and mixture restoration, MDD13 for deblurring evaluation, exBeDDE and IVCDehazing for dehazing assessment, IVIPC-DQA for deraining quality evaluation, and super-resolution datasets QADS, RealSRQ, and SISRSet covering different scenarios. The results consistently demonstrate the same pattern observed in our main analysis: while most methods achieve reasonable overall correlation coefficients when evaluated on complete datasets, their performance dramatically deteriorates within narrow quality ranges. This reveals the inability of traditional methods to effectively conduct fine-grained image quality assessment.
Table 5: Performance analysis across different MOS score ranges on PIPAL dataset. Methods are evaluated on subsets with MOS scores in specific ranges.
Type Method [0.0,0.2) [0.2,0.4) [0.4,0.6) [0.6,0.8) [0.8,1.0] [0.0,1.0]
SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC
FR PSNR 0.323 0.266 0.082 0.107 0.209 0.224 0.161 0.135 0.072 0.021 0.422 0.420
SSIM 0.293 0.243 0.108 0.101 0.258 0.273 0.254 0.195 0.049 0.076 0.530 0.516
LPIPS -0.034 0.025 0.077 0.080 0.325 0.323 0.287 0.267 0.124 0.096 0.612 0.595
DISTS 0.168 0.134 0.159 0.165 0.310 0.304 0.242 0.227 0.165 0.128 0.585 0.584
NR NIQE -0.126 -0.101 -0.002 0.036 0.107 0.114 0.001 -0.096 0.080 0.038 0.153 0.018
IL-NIQE -0.235 -0.144 -0.098 -0.093 0.126 0.128 0.128 0.141 0.054 0.053 0.289 0.274
BRISQUE -0.142 -0.105 0.025 0.014 0.125 0.100 0.035 -0.048 0.131 0.096 0.185 0.078
DB-CNN -0.157 -0.072 0.321 0.316 0.353 0.362 0.330 0.354 -0.016 0.060 0.636 0.669
HyperIQA 0.100 0.174 0.274 0.259 0.314 0.307 0.292 0.298 0.032 -0.045 0.584 0.611
MetaIQA 0.037 0.074 0.160 0.163 0.204 0.196 0.174 0.178 -0.101 -0.060 0.423 0.432
LIQE -0.232 -0.035 0.053 0.059 0.175 0.084 0.299 0.240 0.107 0.182 0.479 0.318
CLIP-IQA -0.152 -0.130 0.211 0.190 0.238 0.236 0.293 0.299 0.071 0.049 0.530 0.542
Q-Align 0.230 0.271 0.301 0.333 0.337 0.324 0.213 0.220 0.178 0.154 0.418 0.410
DeQA-Score 0.568 0.652 0.676 0.672 0.623 0.636 0.516 0.562 0.350 0.339 0.747 0.777
Table 6: Performance comparison across different MOS ranges on TID2013 dataset. FR and NR represent full-reference and no-reference methods, respectively.
Type Method [0.0,0.2) [0.2,0.4) [0.4,0.6) [0.6,0.8) [0.8,1.0] [0.0,1.0]
SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC
FR PSNR 0.064 -0.275 0.164 0.213 0.303 0.330 0.329 0.327 -0.033 -0.126 0.670 0.651
SSIM -0.018 -0.197 0.192 0.262 0.194 0.178 0.380 0.371 0.175 0.113 0.596 0.633
LPIPS -0.318 -0.010 0.308 0.371 0.269 0.252 0.407 0.356 -0.090 -0.045 0.724 0.731
DISTS -0.136 0.052 0.345 0.419 0.268 0.268 0.343 0.350 0.026 0.060 0.688 0.736
NR NIQE -0.191 -0.119 0.221 0.176 0.054 0.076 0.100 0.151 0.014 -0.003 0.313 0.360
IL-NIQE -0.027 0.004 0.236 0.272 0.158 0.157 0.305 0.288 -0.141 -0.068 0.520 0.597
BRISQUE -0.227 -0.246 0.267 0.210 0.091 0.106 0.195 0.196 0.136 0.160 0.370 0.473
DB-CNN 0.636 0.654 0.621 0.591 0.649 0.636 0.496 0.507 0.293 0.303 0.889 0.905
HyperIQA 0.391 0.453 0.587 0.605 0.391 0.347 0.339 0.333 0.000 0.027 0.810 0.839
MetaIQA 0.400 0.502 0.595 0.618 0.563 0.546 0.560 0.526 0.529 0.631 0.885 0.875
LIQE -0.018 -0.140 0.484 0.321 0.468 0.335 0.513 0.508 0.322 0.375 0.869 0.797
CLIP-IQA 0.582 0.567 0.438 0.446 0.467 0.495 0.403 0.413 0.087 0.040 0.813 0.835
Q-Align 0.346 0.464 0.489 0.497 0.510 0.451 0.469 0.472 0.303 0.349 0.867 0.870
DeQA-Score 0.382 0.448 0.566 0.560 0.572 0.566 0.514 0.525 0.467 0.429 0.891 0.898
Table 7: Performance comparison across different MOS ranges on CSIQ dataset. FR and NR represent full-reference and no-reference methods, respectively.
Type Method [0.0,0.2) [0.2,0.4) [0.4,0.6) [0.6,0.8) [0.8,1.0] [0.0,1.0]
SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC
FR PSNR 0.006 0.236 0.536 0.564 0.010 -0.003 0.305 0.411 0.534 0.620 0.808 0.801
SSIM 0.103 -0.035 0.409 0.455 -0.040 0.012 0.378 0.355 0.646 0.475 0.850 0.793
LPIPS 0.333 0.250 0.441 0.461 0.343 0.331 0.620 0.582 0.657 0.568 0.925 0.903
DISTS 0.055 0.037 0.664 0.661 0.288 0.334 0.508 0.509 0.673 0.593 0.933 0.929
NR NIQE 0.661 0.501 0.424 0.351 -0.320 -0.161 0.316 0.346 0.296 0.260 0.619 0.700
IL-NIQE 0.346 0.086 0.446 0.395 0.367 0.318 0.524 0.428 0.216 0.193 0.809 0.787
BRISQUE 0.200 0.215 0.385 0.278 0.076 0.064 0.302 0.301 0.146 0.172 0.559 0.688
DB-CNN 0.624 0.605 0.696 0.657 0.693 0.638 0.526 0.550 0.448 0.483 0.929 0.942
HyperIQA 0.730 0.629 0.472 0.408 0.340 0.419 0.566 0.575 0.316 0.315 0.858 0.891
MetaIQA 0.640 0.605 0.624 0.575 0.552 0.493 0.477 0.417 0.469 0.450 0.888 0.882
LIQE 0.566 0.476 0.579 0.433 0.469 0.489 0.686 0.687 0.675 0.664 0.947 0.888
CLIP-IQA 0.697 0.591 0.306 0.372 0.165 0.203 0.488 0.498 0.352 0.367 0.842 0.868
Q-Align 0.649 0.549 0.576 0.560 0.409 0.444 0.637 0.607 0.341 0.370 0.877 0.904
DeQA-Score 0.830 0.586 0.521 0.491 0.528 0.569 0.546 0.560 0.552 0.551 0.938 0.952
Table 8: Performance comparison across different MOS ranges on LIVE dataset. FR and NR represent full-reference and no-reference methods, respectively.
Type Method [0.0,0.2) [0.2,0.4) [0.4,0.6) [0.6,0.8) [0.8,1.0] [0.0,1.0]
SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC
FR PSNR 0.687 0.627 0.296 0.233 0.525 0.494 0.618 0.628 0.882 0.922 0.952 0.806
SSIM 0.685 0.654 0.298 0.326 0.624 0.637 0.706 0.682 0.894 0.858 0.971 0.885
LPIPS 0.710 0.655 0.263 0.225 0.625 0.641 0.669 0.639 0.891 0.847 0.960 0.845
DISTS 0.710 0.635 0.286 0.215 0.728 0.702 0.703 0.719 0.880 0.879 0.979 0.936
NR NIQE 0.591 0.610 0.151 0.100 0.442 0.449 0.677 0.581 0.474 0.535 0.931 0.645
IL-NIQE 0.677 0.612 0.220 0.113 0.610 0.619 0.501 0.540 0.297 0.196 0.921 0.690
BRISQUE 0.576 0.546 0.506 0.543 0.765 0.768 0.864 0.808 0.541 0.515 0.971 0.976
DB-CNN 0.476 0.486 0.486 0.615 0.797 0.770 0.816 0.780 0.666 0.676 0.977 0.982
HyperIQA 0.676 0.694 0.605 0.715 0.813 0.803 0.856 0.834 0.625 0.655 0.977 0.986
MetaIQA 0.786 0.604 0.581 0.589 0.494 0.462 0.284 0.157 0.266 0.131 0.874 0.837
LIQE -0.185 -0.169 0.461 0.397 0.791 0.769 0.693 0.691 0.365 0.287 0.959 0.943
CLIP-IQA 0.428 0.410 0.452 0.518 0.629 0.615 0.736 0.700 0.500 0.541 0.961 0.968
Q-Align 0.889 0.994 0.907 0.953 0.803 0.716 0.891 0.926 0.554 0.972 0.923 0.928
DeQA-Score 0.845 0.987 0.895 0.922 0.773 0.842 0.895 0.933 0.877 0.991 0.974 0.976
Table 9: Performance analysis across different MOS score ranges on KADID-10K dataset. Methods are evaluated on subsets with MOS scores in specific ranges.
Type Method [0.0,0.2) [0.2,0.4) [0.4,0.6) [0.6,0.8) [0.8,1.0] [0.0,1.0]
SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC
FR PSNR 0.409 0.370 0.212 0.219 0.147 0.126 0.163 0.168 0.162 0.153 0.675 0.554
SSIM 0.442 0.371 0.210 0.176 0.137 0.176 0.180 0.152 0.288 0.293 0.609 0.581
LPIPS 0.416 0.323 0.355 0.298 0.224 0.231 0.276 0.233 0.412 0.399 0.825 0.752
DISTS 0.365 0.315 0.312 0.285 0.244 0.241 0.311 0.301 0.473 0.458 0.807 0.799
NR NIQE 0.036 0.043 0.091 0.076 0.043 0.058 0.061 0.051 0.366 0.337 0.386 0.410
IL-NIQE -0.059 -0.048 0.095 0.111 0.057 0.089 0.160 0.192 0.290 0.302 0.526 0.543
BRISQUE 0.082 0.039 0.045 0.020 0.048 0.037 0.041 0.047 0.256 0.272 0.323 0.387
DB-CNN 0.004 0.009 0.255 0.248 0.311 0.298 0.382 0.386 0.337 0.327 0.822 0.809
HyperIQA 0.047 0.022 0.275 0.256 0.310 0.306 0.373 0.362 0.519 0.460 0.803 0.806
MetaIQA 0.023 0.006 0.159 0.178 0.228 0.208 0.367 0.362 0.487 0.474 0.793 0.782
LIQE 0.194 0.145 0.318 0.284 0.343 0.335 0.359 0.270 0.285 0.195 0.798 0.758
CLIP-IQA 0.125 0.110 0.228 0.235 0.217 0.215 0.313 0.317 0.374 0.346 0.790 0.790
Q-Align 0.436 0.426 0.488 0.482 0.534 0.502 0.534 0.436 0.604 0.445 0.958 0.962
DeQA-Score 0.468 0.532 0.638 0.622 0.673 0.742 0.689 0.762 0.726 0.753 0.926 0.923
Table 10: Performance analysis across different MOS score ranges on VCL@FER. Methods are evaluated on subsets with MOS scores in specific ranges.
Type Method [0.0,0.2) [0.2,0.4) [0.4,0.6) [0.6,0.8) [0.8,1.0] [0.0,1.0]
SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC
FR PSNR 0.114 0.111 0.159 0.178 0.316 0.287 0.422 0.429 0.413 0.488 0.846 0.857
SSIM 0.414 0.480 0.483 0.484 0.398 0.432 0.488 0.381 0.063 0.242 0.919 0.812
LPIPS 0.569 0.606 0.575 0.569 0.449 0.423 0.488 0.358 0.301 0.289 0.896 0.808
DISTS 0.546 0.650 0.630 0.650 0.532 0.562 0.545 0.530 0.462 0.329 0.942 0.913
NR NIQE 0.185 0.089 0.461 0.414 0.413 0.373 0.515 0.439 -0.119 0.174 0.835 0.549
IL-NIQE 0.072 0.100 0.516 0.414 0.306 0.257 0.403 0.385 0.301 0.110 0.796 0.761
BRISQUE 0.274 0.325 0.525 0.409 0.607 0.600 0.448 0.451 0.601 0.571 0.917 0.904
DB-CNN 0.654 0.628 0.733 0.689 0.577 0.627 0.695 0.713 0.273 0.239 0.960 0.959
HyperIQA 0.725 0.711 0.779 0.740 0.558 0.608 0.647 0.612 -0.112 -0.134 0.946 0.945
MetaIQA 0.336 0.384 0.277 0.293 0.233 0.110 0.216 0.169 0.123 -0.115 0.687 0.635
LIQE 0.728 0.400 0.745 0.742 0.630 0.677 0.654 0.612 0.406 0.493 0.963 0.954
CLIP-IQA 0.345 0.291 0.547 0.542 0.356 0.372 0.708 0.696 0.371 0.389 0.952 0.958
Q-Align 0.718 0.275 0.576 0.629 0.435 0.380 0.633 0.692 0.594 0.607 0.890 0.858
DeQA-Score 0.631 0.631 0.720 0.782 0.678 0.717 0.812 0.809 0.559 0.661 0.957 0.956
Table 11: Performance analysis across different MOS score ranges on exBeDDE. Methods are evaluated on subsets with MOS scores in specific ranges.
Type Method [0.0,0.2) [0.2,0.4) [0.4,0.6) [0.6,0.8) [0.8,1.0] [0.0,1.0]
SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC
FR PSNR -0.235 -0.234 0.077 0.049 -0.290 -0.225 -0.304 -0.322 -0.233 -0.309 -0.705 -0.701
SSIM -0.176 -0.177 0.002 0.002 -0.027 0.017 -0.311 -0.320 -0.455 -0.491 -0.730 -0.720
LPIPS -0.075 -0.055 -0.117 -0.189 -0.421 -0.450 -0.245 -0.246 -0.321 -0.403 -0.747 -0.752
DISTS 0.001 -0.060 -0.190 -0.236 -0.322 -0.285 -0.404 -0.341 -0.263 -0.227 -0.783 -0.755
NR NIQE 0.123 0.091 -0.082 -0.065 0.198 0.133 -0.128 -0.145 -0.050 -0.055 0.029 0.031
IL-NIQE 0.108 0.082 -0.036 -0.039 0.032 0.061 0.014 -0.032 -0.402 -0.335 0.107 0.074
BRISQUE -0.086 -0.091 0.116 0.140 0.180 0.195 -0.223 -0.237 0.241 0.220 0.118 0.068
DB-CNN 0.155 0.197 0.344 0.371 0.200 0.264 0.396 0.384 0.731 0.721 0.899 0.905
HyperIQA 0.125 0.191 0.324 0.298 0.226 0.206 0.159 0.143 0.594 0.610 0.847 0.863
MetaIQA 0.135 0.239 0.297 0.298 0.262 0.263 0.223 0.221 0.662 0.617 0.860 0.881
LIQE 0.017 0.143 0.433 0.407 0.383 0.356 0.385 0.376 0.734 0.492 0.895 0.893
CLIP-IQA 0.168 0.179 0.397 0.406 0.328 0.342 0.215 0.217 0.524 0.491 0.879 0.903
Q-Align 0.013 0.044 0.153 0.170 0.218 0.260 0.286 0.289 0.472 0.432 0.814 0.815
DeQA-Score 0.105 0.142 0.201 0.186 0.172 0.275 0.258 0.265 0.707 0.727 0.863 0.879
Table 12: Performance analysis across different MOS score ranges on IVCDehazing. Methods are evaluated on subsets with MOS scores in specific ranges.
Type Method [0.0,0.2) [0.2,0.4) [0.4,0.6) [0.6,0.8) [0.8,1.0] [0.0,1.0]
SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC
FR PSNR - - - - - - - - - - - -
SSIM - - - - - - - - - - - -
LPIPS - - - - - - - - - - - -
DISTS - - - - - - - - - - - -
NR NIQE -1.000 -1.000 -0.113 -0.056 0.127 0.059 0.193 0.150 0.600 0.920 0.357 0.455
IL-NIQE 1.000 1.000 -0.042 -0.162 0.582 0.440 0.132 0.105 -0.200 -0.576 0.381 0.509
BRISQUE 1.000 1.000 0.459 0.407 -0.291 -0.022 0.062 -0.003 0.200 0.582 0.122 0.124
DB-CNN 1.000 1.000 -0.127 -0.175 0.255 0.268 0.324 0.397 -0.200 0.310 0.753 0.654
HyperIQA 1.000 1.000 -0.177 -0.101 0.800 0.715 0.130 0.002 -0.400 -0.667 0.404 0.592
MetaIQA 1.000 1.000 0.127 0.150 -0.346 -0.422 0.171 0.270 0.000 0.085 0.683 0.618
LIQE 1.000 1.000 -0.438 -0.513 0.255 0.455 0.443 0.380 0.200 0.377 0.630 0.540
CLIP-IQA 1.000 1.000 0.223 0.132 -0.255 -0.332 0.107 0.179 0.000 0.154 0.530 0.559
Q-Align 1.000 1.000 0.329 0.278 0.143 0.072 0.600 0.634 -1.000 -0.978 0.328 0.303
DeQA-Score 1.000 1.000 0.410 0.532 0.643 0.778 -0.150 -0.323 -0.400 -0.460 0.103 0.177
Table 13: Performance analysis across different MOS score ranges on IVIPC-DQA dataset. Methods are evaluated on subsets with MOS scores in specific ranges.
Type Method [0.0,0.2) [0.2,0.4) [0.4,0.6) [0.6,0.8) [0.8,1.0] [0.0,1.0]
SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC
FR PSNR - - - - - - - - - - - -
SSIM - - - - - - - - - - - -
LPIPS - - - - - - - - - - - -
DISTS - - - - - - - - - - - -
NR NIQE 0.255 0.240 0.077 0.073 0.121 0.077 - - - - 0.032 0.008
IL-NIQE -0.094 -0.073 0.038 0.045 0.184 0.126 - - - - 0.002 0.033
BRISQUE -0.003 -0.115 -0.018 -0.028 0.113 0.172 - - - - 0.090 0.094
DB-CNN 0.164 0.165 0.338 0.369 0.238 0.264 - - - - 0.491 0.525
HyperIQA 0.011 0.038 0.369 0.417 0.050 0.111 - - - - 0.475 0.503
MetaIQA -0.022 -0.040 0.393 0.353 0.319 0.364 - - - - 0.550 0.546
LIQE -0.098 0.044 0.329 0.332 0.136 0.152 - - - - 0.483 0.479
CLIP-IQA -0.043 -0.041 0.159 0.189 0.215 0.240 - - - - 0.400 0.450
Q-Align 0.015 -0.105 0.272 0.349 0.245 0.324 - - - - 0.255 0.302
DeQA-Score 0.571 0.544 0.512 0.602 0.506 0.502 - - - - 0.533 0.581
Table 14: Performance analysis across different MOS score ranges on MDD13 dataset. Methods are evaluated on subsets with MOS scores in specific ranges.
Type Method [0.0,0.2) [0.2,0.4) [0.4,0.6) [0.6,0.8) [0.8,1.0] [0.0,1.0]
SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC
FR PSNR 1.000 0.656 0.182 0.154 0.321 0.317 0.048 0.066 0.014 -0.068 0.186 0.168
SSIM 1.000 0.657 0.015 -0.008 0.261 0.294 0.289 0.167 0.125 0.071 0.445 0.353
LPIPS 0.800 0.799 0.232 0.093 0.252 0.091 0.340 0.326 0.472 0.451 0.777 0.702
DISTS 1.000 0.922 0.400 0.336 0.295 0.292 0.600 0.591 0.790 0.820 0.907 0.901
NR NIQE -0.200 -0.702 0.296 0.302 0.128 0.097 0.253 0.287 0.372 0.475 0.386 0.413
IL-NIQE 0.000 -0.287 0.331 0.127 0.176 0.150 0.258 0.287 0.394 0.382 0.591 0.515
BRISQUE 0.200 0.029 -0.130 -0.238 0.091 0.075 0.304 0.304 0.206 0.227 0.356 0.356
DB-CNN 0.800 0.698 0.540 0.671 0.437 0.447 0.538 0.551 0.476 0.515 0.868 0.888
HyperIQA 0.000 0.318 0.577 0.603 0.370 0.388 0.551 0.551 0.377 0.387 0.862 0.879
MetaIQA 1.000 0.843 0.455 0.445 0.310 0.279 0.462 0.391 0.282 0.218 0.806 0.826
LIQE 0.949 0.656 0.536 0.559 0.497 0.442 0.594 0.582 0.370 0.400 0.899 0.814
CLIP-IQA 0.316 0.649 0.449 0.610 0.499 0.524 0.604 0.628 0.556 0.622 0.908 0.910
Q-Align -0.400 -0.086 0.380 0.488 0.296 0.301 0.589 0.588 0.672 0.663 0.884 0.880
DeQA-Score 0.800 0.916 0.432 0.608 0.429 0.445 0.529 0.532 0.493 0.539 0.887 0.905
Table 15: Performance analysis across different MOS score ranges on QADS dataset. Methods are evaluated on subsets with MOS scores in specific ranges.
Type Method [0.0,0.2) [0.2,0.4) [0.4,0.6) [0.6,0.8) [0.8,1.0] [0.0,1.0]
SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC
FR PSNR 1.000 0.656 0.182 0.154 0.321 0.317 0.048 0.066 0.014 -0.068 0.186 0.168
SSIM 1.000 0.657 0.015 -0.008 0.261 0.294 0.289 0.167 0.125 0.071 0.445 0.353
LPIPS 0.800 0.799 0.232 0.093 0.252 0.091 0.340 0.326 0.472 0.451 0.777 0.702
DISTS 1.000 0.922 0.400 0.336 0.295 0.292 0.600 0.591 0.790 0.820 0.907 0.901
NR NIQE -0.200 -0.702 0.296 0.302 0.128 0.097 0.253 0.287 0.372 0.475 0.386 0.413
IL-NIQE 0.000 -0.287 0.331 0.127 0.176 0.150 0.258 0.287 0.394 0.382 0.591 0.515
BRISQUE 0.200 0.029 -0.130 -0.238 0.091 0.075 0.304 0.304 0.206 0.227 0.356 0.356
DB-CNN 0.800 0.698 0.540 0.671 0.437 0.447 0.538 0.551 0.476 0.515 0.868 0.888
HyperIQA 0.000 0.318 0.577 0.603 0.370 0.388 0.551 0.551 0.377 0.387 0.862 0.879
MetaIQA 1.000 0.843 0.455 0.445 0.310 0.279 0.462 0.391 0.282 0.218 0.806 0.826
LIQE 0.949 0.656 0.536 0.559 0.497 0.442 0.594 0.582 0.370 0.400 0.899 0.814
CLIP-IQA 0.316 0.649 0.449 0.610 0.499 0.524 0.604 0.628 0.556 0.622 0.908 0.910
Q-Align -0.400 -0.086 0.380 0.488 0.296 0.301 0.589 0.588 0.672 0.663 0.884 0.880
DeQA-Score 0.800 0.916 0.432 0.608 0.429 0.445 0.529 0.532 0.493 0.539 0.887 0.905
Table 16: Performance analysis across different MOS score ranges on RealSRQ dataset. Methods are evaluated on subsets with MOS scores in specific ranges.
Type Method [0.0,0.2) [0.2,0.4) [0.4,0.6) [0.6,0.8) [0.8,1.0] [0.0,1.0]
SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC
FR PSNR 1.000 1.000 -0.149 -0.151 -0.069 0.010 - - - - -0.035 0.013
SSIM 1.000 1.000 -0.173 -0.116 0.021 0.118 - - - - -0.029 0.050
LPIPS 1.000 1.000 -0.148 -0.106 0.271 0.202 - - - - -0.003 0.071
DISTS 1.000 1.000 -0.147 -0.132 0.311 0.309 - - - - -0.034 0.057
NR NIQE -1.000 -1.000 -0.017 -0.005 0.359 0.362 - - - - 0.013 0.050
IL-NIQE -1.000 -1.000 -0.031 -0.014 0.218 0.205 - - - - 0.006 0.032
BRISQUE -1.000 -1.000 0.059 0.117 0.095 0.156 - - - - 0.045 -0.018
DB-CNN -1.000 -1.000 0.610 0.630 0.786 0.801 - - - - 0.739 0.838
HyperIQA -1.000 -1.000 0.474 0.490 0.705 0.743 - - - - 0.602 0.609
MetaIQA 1.000 1.000 0.106 -0.003 -0.061 0.029 - - - - 0.171 0.134
LIQE -1.000 -1.000 0.412 0.347 0.561 0.421 - - - - 0.537 0.525
CLIP-IQA -1.000 -1.000 0.395 0.439 0.636 0.660 - - - - 0.555 0.659
Q-Align -1.000 -1.000 -0.058 0.005 0.201 0.303 - - - - 0.100 0.227
DeQA-Score -1.000 -1.000 0.267 0.269 0.685 0.679 - - - - 0.445 0.609
Table 17: Performance analysis across different MOS score ranges on SISRSet dataset. Methods are evaluated on subsets with MOS scores in specific ranges.
Type Method [0.0,0.2) [0.2,0.4) [0.4,0.6) [0.6,0.8) [0.8,1.0] [0.0,1.0]
SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC
FR PSNR -0.314 -0.168 -0.136 0.077 0.261 0.077 0.762 0.741 -0.367 -0.231 0.589 0.613
SSIM -0.074 -0.072 -0.009 -0.068 0.029 -0.240 0.833 0.648 -0.346 -0.196 0.371 0.406
LPIPS 0.279 0.165 -0.600 -0.511 -0.540 -0.592 -0.167 0.129 -0.217 -0.364 0.517 0.560
DISTS 0.220 0.122 -0.409 -0.427 -0.503 -0.576 -0.048 0.242 -0.045 -0.415 0.665 0.653
NR NIQE 0.288 0.260 0.446 0.525 0.753 0.809 -0.191 -0.238 0.351 0.430 0.680 0.683
IL-NIQE 0.712 0.687 0.236 0.371 0.562 0.594 -0.071 -0.120 0.338 0.449 0.639 0.616
BRISQUE 0.664 0.680 0.055 -0.064 0.213 0.444 -0.310 -0.194 0.512 0.458 0.888 0.876
DB-CNN 0.664 0.655 0.791 0.675 0.852 0.804 0.476 0.040 0.772 0.683 0.915 0.902
HyperIQA 0.801 0.810 0.636 0.642 0.617 0.623 0.262 0.053 0.452 0.491 0.959 0.946
MetaIQA 0.444 0.385 0.091 0.180 0.050 0.435 -0.357 -0.477 0.479 0.222 0.638 0.511
LIQE 0.191 0.266 0.364 0.100 0.631 0.596 -0.333 -0.476 0.397 0.122 0.307 0.436
CLIP-IQA 0.606 0.653 0.164 0.225 0.125 0.043 0.071 0.032 0.436 0.371 0.945 0.941
Q-Align 0.805 0.799 0.564 0.528 0.208 0.197 -0.119 -0.175 -0.204 -0.081 0.679 0.665
DeQA-Score 0.553 0.528 0.427 0.360 0.356 0.336 -0.381 -0.282 0.591 0.191 0.496 0.417
Appendix C Fine-grained Image Pairs Visualization
This section presents visual examples demonstrating the perceptual challenges in fine-grained quality assessment across six restoration tasks. These examples represent the subtle quality variations commonly encountered in practical image restoration applications, where algorithmic comparisons and parameter optimization require sensitive discrimination between restoration results with marginal quality differences.
Figures 7 through 12 showcase image pairs with subtle quality differences in deblurring, denoising, deraining, super-resolution, dehazing, and mixture restoration tasks. The image pairs demonstrate cases where slight variations in restoration effectiveness can significantly impact perceptual quality despite producing similar overall visual appearance.
These carefully selected examples illustrate the inherent complexity of fine-grained quality assessment in image restoration. The visual evidence demonstrates that distinguishing subtle quality differences requires sophisticated understanding of restoration-specific artifacts and perceptual quality factors. These examples highlight the necessity for FGResQ designed to handle the nuanced quality differences encountered in image restoration applications, where traditional evaluation paradigms may not capture the full spectrum of perceptual quality variations.
Figure 7: Fine-grained image pairs for Deblurring
Figure 8: Fine-grained image pairs for Denoising
Figure 9: Fine-grained image pairs for Deraining
Figure 10: Fine-grained image pairs for Super-Resolution
Figure 11: Fine-grained image pairs for Dehazing
Figure 12: Fine-grained image pairs for Mixture Restoration
Appendix D Qualitative Analysis
This section provides visual evidence of how different quality assessment methods respond to fine-grained quality differences in restored images. Figures 13 through 18 present qualitative comparisons on representative fine-grained image pairs across different restoration tasks. The results demonstrate systematic limitations of existing IQA methods in fine-grained scenarios.
Traditional metrics like PSNR and SSIM often produce nearly identical scores for images with subtle quality differences, failing to provide meaningful quality discrimination. Advanced learning-based methods such as CLIP-IQA and DeQA-Score show improved sensitivity compared to traditional metrics but still exhibit inconsistent behavior in fine-grained evaluation. These methods frequently produce incorrect rankings when quality differences become subtle, particularly in challenging scenarios like mixture restoration where multiple degradation types interact.
In contrast, our proposed FGResQ demonstrates superior performance across all tested scenarios. The method consistently identifies the superior image in fine-grained pairs where existing approaches fail, particularly excelling in challenging mixture restoration and dehazing scenarios. FGResQ’s degradation-aware feature learning enables more nuanced understanding of restoration quality, while the dual-branch architecture effectively captures both coarse-grained and fine-grained quality relationships. These visual comparisons complement our quantitative analysis and provide intuitive evidence demonstrating FGResQ’s effectiveness in addressing the fundamental limitations of existing approaches for fine-grained quality assessment in image restoration applications.
Figure 13: Qualitative comparison for Deblurring
Figure 14: Qualitative comparison for Denoising
Figure 15: Qualitative comparison for Deraining
Figure 16: Qualitative comparison for Super-Resolution
Figure 17: Qualitative comparison for Dehazing
Figure 18: Qualitative comparison for Mixture Restoration

### v3

@@
 Extensive experiments and comparisons demonstrate that FGResQ significantly outperforms state-of-the-art IQA metrics.
-Codes and model weights have been released in https://pxf0429.github.io/FGResQ/.
+Codes and model weights have been released in https://sxfly99.github.io/FGResQ-Homepage.
 Introduction
-Perceptual image restoration (IR) stands as a cornerstone in low-level computer vision, aiming to recover high-quality images from their degraded observations while maintaining perceptual fidelity (Potlapalli et al.
-2023; Luo et al.
-2023).
+Perceptual image restoration (IR) stands as a cornerstone in low-level computer vision, aiming to recover high-quality images from their degraded observations while maintaining perceptual fidelity (potlapalli2023promptir; luo2023controlling).
 Recent years have witnessed remarkable breakthroughs in this field, largely driven by the evolution of generative models.
@@
 Such assessment capabilities are essential not only for performance benchmarking across different IR algorithms but also for guiding algorithmic optimization.
-Currently, IR algorithm evaluation still predominantly relies on traditional reference-based metrics such as PSNR and SSIM (Wang et al.
-2004), which assess restoration quality by measuring the similarity between the restored images and ground-truth images, even though their inconsistency with human perceptual judgment has been increasingly observed by the research community (Jinjin et al.
-2020).
-Recent efforts have introduced no-reference IQA methods (Xu et al.
-2024; Zhou et al.
-2025), including CLIP-IQA (Wang, Chan, and Loy 2023) and other learning-based approaches (Ke et al.
-2021; Wu et al.
-2023), which prove particularly valuable in real-world restoration scenarios where ground truth images are unavailable.
-However, a critical oversight persists: these methods have been primarily designed for image quality assessment, without specific consideration for the fine-grained evaluation requirements inherent in IR tasks.
 Figure 1: Illustration of the fine-grained challenge in IQA for image restoration.
@@
 (b) FGResQ enables both coarse-grained quality scoring and fine-grained quality ranking capabilities.
+Currently, IR algorithm evaluation still predominantly relies on traditional reference-based metrics such as PSNR and SSIM (SSIM), which assess restoration quality by measuring the similarity between the restored images and ground-truth images, even though their inconsistency with human perceptual judgment has been increasingly observed by the research community (PIPAL).
+Recent efforts have introduced no-reference IQA methods (xu2024towards; zhou2025unires), including CLIP-IQA (CLIP-IQA) and other learning-based approaches (MUSIQ; Q-align), which prove particularly valuable in real-world restoration scenarios where ground truth images are unavailable.
+However, a critical oversight persists: these methods have been primarily designed for image quality assessment, without specific consideration for the fine-grained evaluation requirements inherent in IR tasks.
 A critical yet underexplored challenge in IR evaluation lies in its inherently fine-grained nature.
@@
 Can existing score-based IQA methods objectively capture fine-grained quality differences in restored images?
-To investigate this question, we first conducted comprehensive computational analyses on established IQA datasets for image restoration (Jinjin et al.
-2020; Shi et al.
-2019).
+To investigate this question, we first conducted comprehensive computational analyses on established IQA datasets for image restoration (PIPAL; SISRSet).
 Our findings consistently reveal a significant performance gap: while existing IQA methods achieve reasonable results in coarse-grained quality assessment, they consistently underperform in fine-grained scenarios, with significant inconsistencies between predicted scores and human perceptual judgments.
@@
 Moreover, when quality differences between images become subtle, both human and IQA models encounter significant challenges in providing reliable absolute quality scores.
-In contrast, pairwise comparison presents a more reliable alternative for fine-grained assessment, as humans demonstrate superior consistency in relative quality judgments compared to absolute scoring.
+In contrast, pairwise comparison(10168279; yang2024semantics) presents a more reliable alternative for fine-grained assessment, as humans demonstrate superior consistency in relative quality judgments compared to absolute scoring.
 Motivated by the above insights, we introduce FGRestore, the first fine-grained IQA dataset specifically for image restoration tasks, as shown in Figure 2(a).
@@
 To evaluate IR algorithms, numerous datasets have been proposed.
-These include datasets dedicated to single restoration types, such as the MDD13 (Liu et al.
-2013) dataset for deblurring, the IVC-Dehazing (Ma, Liu, and Wang 2015) and exBeDDE (Zhao et al.
-2020) datasets for dehazing, the IVIPC-DQA (Wu et al.
-2019) dataset for deraining, and the QADS (Zhou et al.
-2019), SISRSet (Shi et al.
-2019), RealSRQ (Jiang et al.
-2022) and SRIQA-Bench (Chen et al.
-2025) datasets for super-resolution.
-Additionally, there are datasets that encompass distortions from various algorithms, such as PIPAL (Jinjin et al.
-2020), which includes distortion types like denoising, super-resolution, and super-resolution with post-denoising, KADID-10k (Lin, Hosu, and Saupe 2019), which contains various image distortions including denoising degradations and DiffIQA(Chen et al.
-2025), which accounts for degradation artifacts induced by diffusion-based image enhancement.
+These include datasets dedicated to single restoration types, such as the MDD13 (MDD13) dataset for deblurring, the IVC-Dehazing (IVC-Dehazing) and exBeDDE (exBeDDE) datasets for dehazing, the IVIPC-DQA (IVIPC_DQA) dataset for deraining, and the QADS (QADS), SISRSet (SISRSet), RealSRQ (RealSRQ) and SRIQA-Bench (chen2025toward) datasets for super-resolution.
+Additionally, there are datasets that encompass distortions from various algorithms, such as PIPAL (PIPAL), which includes distortion types like denoising, super-resolution, and super-resolution with post-denoising, KADID-10k (KADID-10k), which contains various image distortions including denoising degradations and DiffIQA(chen2025toward), which accounts for degradation artifacts induced by diffusion-based image enhancement.
 A brief comparison of FGRestore and previous IR evaluation datasets is illustrated in Table 1.
-Image Quality Assessment
-Existing image restoration (IR) algorithms typically employ full-reference image quality assessment (FR-IQA) metrics for performance evaluation, such as the conventional PSNR and SSIM (Wang et al.
-2004), as well as deep learning-based metrics like LPIPS (Zhang et al.
-2018a) and DISTS (Ding et al.
-2020).
-To overcome the flawed ”perfect reference” assumption in traditional FR-IQA, which is challenged by imaging system limitations and superior generative methods, A-FINE (Chen et al.
-2025) proposes a generalized model that adaptively assesses both the fidelity and naturalness of test images.
-No-reference IQA (NR-IQA) methods have been proposed to assess image quality without reference images.
-Early NR-IQA methods, such as NIQE (Mittal, Soundararajan, and Bovik 2012), primarily rely on hand-crafted Natural Scene Statistics (NSS) features.
-With the advancement of deep learning, many works have turned to neural networks for NR-IQA.
-For instance, MetaIQA (Zhu et al.
-2020) leverages meta-learning to enhance model generalization.
-Additionally, methods such as CLIP-IQA (Wang, Chan, and Loy 2023) leverage text prompts to enhance perceptual capabilities.
-Given the powerful learning capabilities of large multi-modality models (LMMs), emerging work, such as Q-Align (Wu et al.
-2023), has begun to apply these models to quality assessment.
-Most recently, methods represented by Co-Instruct (Wu et al.
-2024) have extended LMMs to compare multiple images.
-Furthermore, VisualQuality-R1(Wu et al.
-2025) introduces a novel methodology for IQA that leverages LMMs within a reinforcement learning-to-rank framework.
-Although these datasets and methods have contributed significantly to the IQA community, most methods are score-based and cannot effectively handle fine-grained perceptual differences brought by image restoration algorithms.
-Recent works have explored fine-grained quality assessment in specific domains.
-Zhang et al.
-(Zhang et al.
-2018b; Zhang, Lin, and Huang 2021) introduced fine-grained assessment for compressed images, focusing on distinguishing compression artifacts within similar quality levels.
-However, these prior works primarily target image compression, whereas our work specifically addresses the unique challenges of fine-grained quality assessment in image restoration tasks, where the evaluation focus shifts from compression artifacts to perceptual restoration quality differences.
-Table 1: Comparison with the previous datasets.
-MOS: Mean Opinion Score; PC: pairwise comparison; ACR: Absolute Category Rating; DCR: Degradation Category Rating; 5-level: 5-point quality scale.
-Number show Images/Pairs/Annotations.
 Dataset Year Task Alg. Annotation Number
@@
 FGRestore 2025 6 108 MOS+RANK 18,408/30,886/45,318
-Table 2: Performance comparison across different MOS ranges on PIPAL dataset.
+Table 1: Comparison with the previous datasets.
+MOS: Mean Opinion Score; PC: pairwise comparison; ACR: Absolute Category Rating; DCR: Degradation Category Rating; 5-level: 5-point quality scale.
+Number show Images/Pairs/Annotations.
+Image Quality Assessment
+Existing image restoration (IR) algorithms typically employ full-reference image quality assessment (FR-IQA) metrics for performance evaluation, such as the conventional PSNR and SSIM (SSIM), as well as deep learning-based metrics like LPIPS (LPIPS) and DISTS (DISTS).
+To overcome the flawed ”perfect reference” assumption in traditional FR-IQA, which is challenged by imaging system limitations and superior generative methods, A-FINE (chen2025toward) proposes a generalized model that adaptively assesses both the fidelity and naturalness of test images.
+No-reference IQA (NR-IQA) methods have been proposed to assess image quality without reference images.
+Early NR-IQA methods, such as NIQE(NIQE), primarily rely on hand-crafted Natural Scene Statistics (NSS) features.
+With the advancement of deep learning, many works have turned to neural networks for NR-IQA.
+For instance, MetaIQA(MetaIQA)leverages meta-learning to enhance model generalization.
+Additionally, methods such as CLIP-IQA(CLIP-IQA)leverage text prompts to enhance perceptual capabilities(sheng2023aesclip; yang2025language; li2025cgiaa).
+Given the powerful learning capabilities of large multi-modality models (LMMs)(sheng2025instructcrop; sheng2025agiqa), emerging work, such as Q-Align(Q-align), has begun to apply these models to quality assessment.
+Most recently, methods represented by Co-Instruct(Co-Instruct)have extended LMMs to compare multiple images.
+Furthermore, VisualQuality-R1(wu2025visualquality) introduces a novel methodology for IQA that leverages LMMs within a reinforcement learning-to-rank framework.
+Although these datasets and methods have contributed significantly to the IQA community, most methods are score-based and cannot effectively handle fine-grained perceptual differences brought by image restoration algorithms.
+Recent works have explored fine-grained quality assessment in specific domains.
+Zhang et al.
+(zhang2019fine; zhang2022fine) introduced fine-grained assessment for compressed images, focusing on distinguishing compression artifacts within similar quality levels.
+However, these prior works primarily target image compression, whereas our work specifically addresses the unique challenges of fine-grained quality assessment in image restoration tasks, where the evaluation focus shifts from compression artifacts to perceptual restoration quality differences.
 Type Method [0.0,0.2) [0.2,0.4) [0.4,0.6) [0.6,0.8) [0.8,1.0] Overall
@@
 DeQA-Score 0.568 0.676 0.623 0.516 0.350 0.747
+Table 2: Performance comparison across different MOS ranges on PIPAL dataset.
 Figure 3: Overview of the FGRestore dataset construction.
@@
 Our evaluation encompasses both full-reference (FR) and no-reference (NR) methods.
-FR methods include PSNR, SSIM (Wang et al.
-2004), LPIPS (Zhang et al.
-2018a), and DISTS (Ding et al.
-2020).
-NR methods include traditional approaches such as NIQE (Mittal, Soundararajan, and Bovik 2012), IL-NIQE (Zhang, Zhang, and Bovik 2015), and BRISQUE (Mittal, Moorthy, and Bovik 2012), as well as deep learning-based methods including DB-CNN (Zhang et al.
-2020), HyperIQA (Su et al.
-2020), MetaIQA (Zhu et al.
-2020), LIQE (Zhang et al.
-2023), CLIP-IQA (Wang, Chan, and Loy 2023), Q-Align (Wu et al.
-2023), and DeQA-Score (You et al.
-2025).
+FR methods include PSNR, SSIM (SSIM), LPIPS (LPIPS), and DISTS (DISTS).
+NR methods include traditional approaches such as NIQE (NIQE), IL-NIQE (IL-NIQE), and BRISQUE (BRISQUE), as well as deep learning-based methods including DB-CNN (DBCNN), HyperIQA (HyperIQA), MetaIQA (MetaIQA), LIQE (LIQE), CLIP-IQA (CLIP-IQA), Q-Align (Q-align), and DeQA-Score (DeQA).
 Specifically, we partition the quality score range into several intervals to examine how IQA methods perform when evaluating images with similar quality levels, which is particularly relevant for fine-grained assessment scenarios in image restoration.
-Table 2 presents the performance comparison across different MOS ranges on the PIPAL dataset (Jinjin et al.
-2020).
+Table 2 presents the performance comparison across different MOS ranges on the PIPAL dataset (PIPAL).
 While most IQA methods achieve reasonable overall SRCC values on the whole dataset, their performance dramatically deteriorates when evaluated within narrow quality ranges.
@@
 Image and Pair Collection
-We collect images from multiple restoration-specific datasets to cover different degradation and visual appearances: 1) deraining images from IVIPC-DQA (Wu et al.
-2019); 2) deblurring images from MDD13 (Liu et al.
-2013); 3) dehazing images from IVC-Dehazing (Ma, Liu, and Wang 2015) and exBeDDE (Zhao et al.
-2020); 4) denoising images from PIPAL (Jinjin et al.
-2020) and KADID-10k (Lin, Hosu, and Saupe 2019); 5) super-resolution images from QADS (Zhou et al.
-2019), RealSRQ (Jiang et al.
-2022), SISRSet (Shi et al.
-2019), and PIPAL (Jinjin et al.
-2020); 6) mixture restoration images from PIPAL (Jinjin et al.
-2020).
+We collect images from multiple restoration-specific datasets to cover different degradation and visual appearances: 1) deraining images from IVIPC-DQA (IVIPC_DQA); 2) deblurring images from MDD13 (MDD13); 3) dehazing images from IVC-Dehazing (IVC-Dehazing) and exBeDDE (exBeDDE); 4) denoising images from PIPAL (PIPAL) and KADID-10k (KADID-10k); 5) super-resolution images from QADS (QADS), RealSRQ (RealSRQ), SISRSet (SISRSet), and PIPAL (PIPAL); 6) mixture restoration images from PIPAL (PIPAL).
 After image collection, we generate image pairs by randomly pairing all images with identical content and restoration tasks, resulting in a total of 1,275,297 image pairs.
@@
 The JND represents the minimum perceptual threshold for noticeable quality differences, making it ideal for identifying imperceptible quality pairs.
-Our JND-based filtering process operates as follows: (1) randomly select 20,000 images from the collected datasets, (2) compute the JND map for each image using established JND estimation model (Wu et al.
-2017), and (3) generate JND noise images by adding these JND maps as noise to the original images:
+Our JND-based filtering process operates as follows: (1) randomly select 20,000 images from the collected datasets, (2) compute the JND map for each image using established JND estimation model (wu2017enhanced), and (3) generate JND noise images by adding these JND maps as noise to the original images:
 $I_{\text{JND}}=I+\text{JND}(I),$ (2)
 where $I$ is the original image and $\text{JND}(I)$ is the corresponding JND noise map.
-We calculate SSIM (Wang et al.
-2004) between each $I$ and $I_{\text{JND}}$, and use the median SSIM value $\text{SSIM}_{\text{med}}$ as our filtering threshold.
+We calculate SSIM (SSIM) between each $I$ and $I_{\text{JND}}$, and use the median SSIM value $\text{SSIM}_{\text{med}}$ as our filtering threshold.
 The unnoticeable pair filtering criterion is then defined as:
@@
 Since FGRestore preserves original score annotations from different source datasets with varying quality scales, we employ a scene-aware fidelity loss that focuses on ranking relationships within each dataset (scene) rather than absolute score values:
-$\mathcal{L}_{fid}^{(s)}$ $=\frac{1}{N_{s}(N_{s}-1)/2}\sum_{i<j}\Big{(}1-\sqrt{p_{ij}g_{ij}+\epsilon}$
-$\quad-\sqrt{(1-p_{ij})(1-g_{ij})+\epsilon}\Big{)},$ (6)
+$\mathcal{L}_{fid}^{(s)}=\frac{1}{\binom{N_{s}}{2}}\sum_{i<j}\Big($ $1-\sqrt{p_{ij}g_{ij}+\epsilon}$
+$-\sqrt{(1-p_{ij})(1-g_{ij})+\epsilon}\Big),$ (6)
 where $p_{ij}=\sigma(y_{pred,i}-y_{pred,j})$, $g_{ij}=\frac{1}{2}(\text{sign}(y_{gt,i}-y_{gt,j})+1)$, and $N_{s}$ is the number of samples in scene $s$.
@@
 where $\lambda_{1}$ and $\lambda_{2}$ are balancing hyperparameters.
-Table 3: Performance comparison on FGRestore dataset across different IR tasks. ”-” indicates no reference images available.
 Type Method Pub. Deblurring Denoising Deraining
@@
 FGResQ 0.821 0.854 0.669 0.698 0.706 0.713 0.521 0.536 0.721 0.703 0.717 0.752
+Table 3: Performance comparison on FGRestore dataset across different IR tasks. ”-” indicates no reference images available.
 Experiments
@@
 We compare the proposed FGResQ model against state-of-the-art FR and NR IQA methods.
-For fair comparison, all learning-based methods with public source codes are retrained using the same training-testing protocol with our scene-aware loss replacing their original regression loss, except for Compare2Score (Zhu et al.
-2024) which operates in a zero-shot manner without training.
+For fair comparison, all learning-based methods with public source codes are retrained using the same training-testing protocol with our scene-aware loss replacing their original regression loss, except for Compare2Score (Compare2Score) which operates in a zero-shot manner without training.
 We evaluate performance using SRCC and PLCC for quality score prediction, and accuracy (ACC) for fine-grained pairwise ranking.
@@
 Implementation Details
-We adopt CLIP ViT-B/16 as the backbone encoder for both degradation-aware feature learning and dual-branch quality prediction.
-Training images are resized to $256\times 256$ pixels and then randomly cropped to $224\times 224$ patches.
-The model is trained using the Adam optimizer with cosine annealing scheduling, where the maximum learning rate is set to $5\times 10^{-6}$, batch size to 64.
+We use CLIP ViT-B/16 as the backbone encoder for both degradation-aware feature learning and dual-branch quality prediction.
+Training images are resized to $256\times 256$ pixels and randomly cropped to $224\times 224$ patches.
+The model is trained on NVIDIA RTX4090 GPUs using the Adam optimizer with cosine annealing.
+The maximum learning rate is $5\times 10^{-6}$ and the batch size is 64.
 To ensure training stability, we employ scene-aware sampling strategy that groups samples from the same source dataset within each batch.
 $\lambda_{1}$ and $\lambda_{2}$ are set to 5 and 1, respectively.
-The model is trained for 6 epochs using all image pairs from the training set on NVIDIA RTX4090 GPUs.
 Performance Evaluation
@@
 This work establishes a new evaluation framework for image restoration and provides valuable insights for developing more perceptually-aligned quality assessment methods in the era of advanced generative restoration models.
-Supplementary Material for Fine-grained Image Quality Assessment for Perceptual Image Restoration
+Acknowledgments
+This work is supported by National Natural Science Foundation of China under Grants 62471349, 62171340 and 62301378, Fundamental Research Funds for the Central Universities under Grant QTZX25076, and partly supported by the China Postdoctoral Science Foundation under Grant 2024M762553.
+Supplementary Material for
+Fine-grained Image Quality Assessment for Perceptual Image Restoration
 This supplementary document provides comprehensive experimental evidence, detailed analyses, and extensive visual demonstrations that substantiate the findings presented in our main paper.
@@
 All source code and implementation details are provided in the supplementary material.
-Appendix A Ablation Study
+Appendix S1 Ablation Study
 As shown in Table 3 in main paper, , unlike task-specific methods that excel in particular scenarios, FGResQ maintains robust performance across all six IR tasks, suggesting that incorporating degradation knowledge enables more generalizable quality assessment capabilities.To further validate the effectiveness of our degradation-aware feature learning (DFL) module, we conduct ablation experiments.
-Table 4 presents the average performance across all IR tasks, comparing our full FGResQ model with a variant that removes the degradation-aware learning component (w/o DFL).
-Table 4: Ablation study.
+Table S1 presents the average performance across all IR tasks, comparing our full FGResQ model with a variant that removes the degradation-aware learning component (w/o DFL).
+Table S1: Ablation study.
 Method SRCC PLCC ACC
@@
 FGResQ 0.703 0.717 0.752
-Appendix B Comprehensive Performance Analysis Across Quality Ranges
+Appendix S2 Comprehensive Performance Analysis Across Quality Ranges
 Following our methodology in the main paper, we partition quality score ranges into discrete intervals, and systematically evaluate existing IQA methods across 13 benchmark datasets.
@@
 This reveals the inability of traditional methods to effectively conduct fine-grained image quality assessment.
-Table 5: Performance analysis across different MOS score ranges on PIPAL dataset. Methods are evaluated on subsets with MOS scores in specific ranges.
+Table S2: Performance analysis across different MOS score ranges on PIPAL dataset. Methods are evaluated on subsets with MOS scores in specific ranges.
 Type Method [0.0,0.2) [0.2,0.4) [0.4,0.6) [0.6,0.8) [0.8,1.0] [0.0,1.0]
@@
 DeQA-Score 0.568 0.652 0.676 0.672 0.623 0.636 0.516 0.562 0.350 0.339 0.747 0.777
-Table 6: Performance comparison across different MOS ranges on TID2013 dataset. FR and NR represent full-reference and no-reference methods, respectively.
+Table S3: Performance comparison across different MOS ranges on TID2013 dataset. FR and NR represent full-reference and no-reference methods, respectively.
 Type Method [0.0,0.2) [0.2,0.4) [0.4,0.6) [0.6,0.8) [0.8,1.0] [0.0,1.0]
@@
 DeQA-Score 0.382 0.448 0.566 0.560 0.572 0.566 0.514 0.525 0.467 0.429 0.891 0.898
-Table 7: Performance comparison across different MOS ranges on CSIQ dataset. FR and NR represent full-reference and no-reference methods, respectively.
+Table S4: Performance comparison across different MOS ranges on CSIQ dataset. FR and NR represent full-reference and no-reference methods, respectively.
 Type Method [0.0,0.2) [0.2,0.4) [0.4,0.6) [0.6,0.8) [0.8,1.0] [0.0,1.0]
@@
 DeQA-Score 0.830 0.586 0.521 0.491 0.528 0.569 0.546 0.560 0.552 0.551 0.938 0.952
-Table 8: Performance comparison across different MOS ranges on LIVE dataset. FR and NR represent full-reference and no-reference methods, respectively.
+Table S5: Performance comparison across different MOS ranges on LIVE dataset. FR and NR represent full-reference and no-reference methods, respectively.
 Type Method [0.0,0.2) [0.2,0.4) [0.4,0.6) [0.6,0.8) [0.8,1.0] [0.0,1.0]
@@
 DeQA-Score 0.845 0.987 0.895 0.922 0.773 0.842 0.895 0.933 0.877 0.991 0.974 0.976
-Table 9: Performance analysis across different MOS score ranges on KADID-10K dataset. Methods are evaluated on subsets with MOS scores in specific ranges.
+Table S6: Performance analysis across different MOS score ranges on KADID-10K dataset. Methods are evaluated on subsets with MOS scores in specific ranges.
 Type Method [0.0,0.2) [0.2,0.4) [0.4,0.6) [0.6,0.8) [0.8,1.0] [0.0,1.0]
@@
 DeQA-Score 0.468 0.532 0.638 0.622 0.673 0.742 0.689 0.762 0.726 0.753 0.926 0.923
-Table 10: Performance analysis across different MOS score ranges on VCL@FER. Methods are evaluated on subsets with MOS scores in specific ranges.
+Table S7: Performance analysis across different MOS score ranges on VCL@FER. Methods are evaluated on subsets with MOS scores in specific ranges.
 Type Method [0.0,0.2) [0.2,0.4) [0.4,0.6) [0.6,0.8) [0.8,1.0] [0.0,1.0]
@@
 DeQA-Score 0.631 0.631 0.720 0.782 0.678 0.717 0.812 0.809 0.559 0.661 0.957 0.956
-Table 11: Performance analysis across different MOS score ranges on exBeDDE. Methods are evaluated on subsets with MOS scores in specific ranges.
+Table S8: Performance analysis across different MOS score ranges on exBeDDE. Methods are evaluated on subsets with MOS scores in specific ranges.
 Type Method [0.0,0.2) [0.2,0.4) [0.4,0.6) [0.6,0.8) [0.8,1.0] [0.0,1.0]
@@
 DeQA-Score 0.105 0.142 0.201 0.186 0.172 0.275 0.258 0.265 0.707 0.727 0.863 0.879
-Table 12: Performance analysis across different MOS score ranges on IVCDehazing. Methods are evaluated on subsets with MOS scores in specific ranges.
+Table S9: Performance analysis across different MOS score ranges on IVCDehazing. Methods are evaluated on subsets with MOS scores in specific ranges.
 Type Method [0.0,0.2) [0.2,0.4) [0.4,0.6) [0.6,0.8) [0.8,1.0] [0.0,1.0]
@@
 DeQA-Score 1.000 1.000 0.410 0.532 0.643 0.778 -0.150 -0.323 -0.400 -0.460 0.103 0.177
-Table 13: Performance analysis across different MOS score ranges on IVIPC-DQA dataset. Methods are evaluated on subsets with MOS scores in specific ranges.
+Table S10: Performance analysis across different MOS score ranges on IVIPC-DQA dataset. Methods are evaluated on subsets with MOS scores in specific ranges.
 Type Method [0.0,0.2) [0.2,0.4) [0.4,0.6) [0.6,0.8) [0.8,1.0] [0.0,1.0]
@@
 DeQA-Score 0.571 0.544 0.512 0.602 0.506 0.502 - - - - 0.533 0.581
-Table 14: Performance analysis across different MOS score ranges on MDD13 dataset. Methods are evaluated on subsets with MOS scores in specific ranges.
+Table S11: Performance analysis across different MOS score ranges on MDD13 dataset. Methods are evaluated on subsets with MOS scores in specific ranges.
 Type Method [0.0,0.2) [0.2,0.4) [0.4,0.6) [0.6,0.8) [0.8,1.0] [0.0,1.0]
@@
 DeQA-Score 0.800 0.916 0.432 0.608 0.429 0.445 0.529 0.532 0.493 0.539 0.887 0.905
-Table 15: Performance analysis across different MOS score ranges on QADS dataset. Methods are evaluated on subsets with MOS scores in specific ranges.
+Table S12: Performance analysis across different MOS score ranges on QADS dataset. Methods are evaluated on subsets with MOS scores in specific ranges.
 Type Method [0.0,0.2) [0.2,0.4) [0.4,0.6) [0.6,0.8) [0.8,1.0] [0.0,1.0]
@@
 DeQA-Score 0.800 0.916 0.432 0.608 0.429 0.445 0.529 0.532 0.493 0.539 0.887 0.905
-Table 16: Performance analysis across different MOS score ranges on RealSRQ dataset. Methods are evaluated on subsets with MOS scores in specific ranges.
+Table S13: Performance analysis across different MOS score ranges on RealSRQ dataset. Methods are evaluated on subsets with MOS scores in specific ranges.
 Type Method [0.0,0.2) [0.2,0.4) [0.4,0.6) [0.6,0.8) [0.8,1.0] [0.0,1.0]
@@
 DeQA-Score -1.000 -1.000 0.267 0.269 0.685 0.679 - - - - 0.445 0.609
-Table 17: Performance analysis across different MOS score ranges on SISRSet dataset. Methods are evaluated on subsets with MOS scores in specific ranges.
+Table S14: Performance analysis across different MOS score ranges on SISRSet dataset. Methods are evaluated on subsets with MOS scores in specific ranges.
 Type Method [0.0,0.2) [0.2,0.4) [0.4,0.6) [0.6,0.8) [0.8,1.0] [0.0,1.0]
@@
 DeQA-Score 0.553 0.528 0.427 0.360 0.356 0.336 -0.381 -0.282 0.591 0.191 0.496 0.417
-Appendix C Fine-grained Image Pairs Visualization
+Appendix S3 Fine-grained Image Pairs Visualization
 This section presents visual examples demonstrating the perceptual challenges in fine-grained quality assessment across six restoration tasks.
 These examples represent the subtle quality variations commonly encountered in practical image restoration applications, where algorithmic comparisons and parameter optimization require sensitive discrimination between restoration results with marginal quality differences.
-Figures 7 through 12 showcase image pairs with subtle quality differences in deblurring, denoising, deraining, super-resolution, dehazing, and mixture restoration tasks.
+Figures S1 through S6 showcase image pairs with subtle quality differences in deblurring, denoising, deraining, super-resolution, dehazing, and mixture restoration tasks.
 The image pairs demonstrate cases where slight variations in restoration effectiveness can significantly impact perceptual quality despite producing similar overall visual appearance.
@@
 These examples highlight the necessity for FGResQ designed to handle the nuanced quality differences encountered in image restoration applications, where traditional evaluation paradigms may not capture the full spectrum of perceptual quality variations.
-Figure 7: Fine-grained image pairs for Deblurring
-Figure 8: Fine-grained image pairs for Denoising
-Figure 9: Fine-grained image pairs for Deraining
-Figure 10: Fine-grained image pairs for Super-Resolution
-Figure 11: Fine-grained image pairs for Dehazing
-Figure 12: Fine-grained image pairs for Mixture Restoration
-Appendix D Qualitative Analysis
+Figure S1: Fine-grained image pairs for Deblurring
+Figure S2: Fine-grained image pairs for Denoising
+Figure S3: Fine-grained image pairs for Deraining
+Figure S4: Fine-grained image pairs for Super-Resolution
+Figure S5: Fine-grained image pairs for Dehazing
+Figure S6: Fine-grained image pairs for Mixture Restoration
+Appendix S4 Qualitative Analysis
 This section provides visual evidence of how different quality assessment methods respond to fine-grained quality differences in restored images.
-Figures 13 through 18 present qualitative comparisons on representative fine-grained image pairs across different restoration tasks.
+Figures S7 through S12 present qualitative comparisons on representative fine-grained image pairs across different restoration tasks.
 The results demonstrate systematic limitations of existing IQA methods in fine-grained scenarios.
@@
 These visual comparisons complement our quantitative analysis and provide intuitive evidence demonstrating FGResQ’s effectiveness in addressing the fundamental limitations of existing approaches for fine-grained quality assessment in image restoration applications.
-Figure 13: Qualitative comparison for Deblurring
-Figure 14: Qualitative comparison for Denoising
-Figure 15: Qualitative comparison for Deraining
-Figure 16: Qualitative comparison for Super-Resolution
-Figure 17: Qualitative comparison for Dehazing
-Figure 18: Qualitative comparison for Mixture Restoration
+Figure S7: Qualitative comparison for Deblurring
+Figure S8: Qualitative comparison for Denoising
+Figure S9: Qualitative comparison for Deraining
+Figure S10: Qualitative comparison for Super-Resolution
+Figure S11: Qualitative comparison for Dehazing
+Figure S12: Qualitative comparison for Mixture Restoration

### v4

@@
 Extensive experiments and comparisons demonstrate that FGResQ significantly outperforms state-of-the-art IQA metrics.
-Codes and model weights have been released in https://sxfly99.github.io/FGResQ-Homepage.
+Homepagehttps://sxfly99.github.io/FGResQ-Home/
 Introduction
-Perceptual image restoration (IR) stands as a cornerstone in low-level computer vision, aiming to recover high-quality images from their degraded observations while maintaining perceptual fidelity (potlapalli2023promptir; luo2023controlling).
+Perceptual image restoration (IR) stands as a cornerstone in low-level computer vision, aiming to recover high-quality images from their degraded observations while maintaining perceptual fidelity (Potlapalli et al.
+2023; Luo et al.
+2023).
 Recent years have witnessed remarkable breakthroughs in this field, largely driven by the evolution of generative models.
@@
 (b) FGResQ enables both coarse-grained quality scoring and fine-grained quality ranking capabilities.
-Currently, IR algorithm evaluation still predominantly relies on traditional reference-based metrics such as PSNR and SSIM (SSIM), which assess restoration quality by measuring the similarity between the restored images and ground-truth images, even though their inconsistency with human perceptual judgment has been increasingly observed by the research community (PIPAL).
-Recent efforts have introduced no-reference IQA methods (xu2024towards; zhou2025unires), including CLIP-IQA (CLIP-IQA) and other learning-based approaches (MUSIQ; Q-align), which prove particularly valuable in real-world restoration scenarios where ground truth images are unavailable.
+Currently, IR algorithm evaluation still predominantly relies on traditional reference-based metrics such as PSNR and SSIM (Wang et al.
+2004), which assess restoration quality by measuring the similarity between the restored images and ground-truth images, even though their inconsistency with human perceptual judgment has been increasingly observed by the research community (Jinjin et al.
+2020).
+Recent efforts have introduced no-reference IQA methods (Xu et al.
+2024; Zhou et al.
+2025), including CLIP-IQA (Wang et al.
+2023) and other learning-based approaches (Ke et al.
+2021), which prove particularly valuable in real-world restoration scenarios where ground truth images are unavailable.
 However, a critical oversight persists: these methods have been primarily designed for image quality assessment, without specific consideration for the fine-grained evaluation requirements inherent in IR tasks.
@@
 Can existing score-based IQA methods objectively capture fine-grained quality differences in restored images?
-To investigate this question, we first conducted comprehensive computational analyses on established IQA datasets for image restoration (PIPAL; SISRSet).
+To investigate this question, we first conducted comprehensive computational analyses on established IQA datasets for image restoration (Jinjin et al.
+2020; Shi et al.
+2019).
 Our findings consistently reveal a significant performance gap: while existing IQA methods achieve reasonable results in coarse-grained quality assessment, they consistently underperform in fine-grained scenarios, with significant inconsistencies between predicted scores and human perceptual judgments.
@@
 Moreover, when quality differences between images become subtle, both human and IQA models encounter significant challenges in providing reliable absolute quality scores.
-In contrast, pairwise comparison(10168279; yang2024semantics) presents a more reliable alternative for fine-grained assessment, as humans demonstrate superior consistency in relative quality judgments compared to absolute scoring.
+In contrast, pairwise comparison(Yang et al.
+2024b, a) presents a more reliable alternative for fine-grained assessment, as humans demonstrate superior consistency in relative quality judgments compared to absolute scoring.
 Motivated by the above insights, we introduce FGRestore, the first fine-grained IQA dataset specifically for image restoration tasks, as shown in Figure 2(a).
@@
 To evaluate IR algorithms, numerous datasets have been proposed.
-These include datasets dedicated to single restoration types, such as the MDD13 (MDD13) dataset for deblurring, the IVC-Dehazing (IVC-Dehazing) and exBeDDE (exBeDDE) datasets for dehazing, the IVIPC-DQA (IVIPC_DQA) dataset for deraining, and the QADS (QADS), SISRSet (SISRSet), RealSRQ (RealSRQ) and SRIQA-Bench (chen2025toward) datasets for super-resolution.
-Additionally, there are datasets that encompass distortions from various algorithms, such as PIPAL (PIPAL), which includes distortion types like denoising, super-resolution, and super-resolution with post-denoising, KADID-10k (KADID-10k), which contains various image distortions including denoising degradations and DiffIQA(chen2025toward), which accounts for degradation artifacts induced by diffusion-based image enhancement.
+These include datasets dedicated to single restoration types, such as the MDD13 (Liu et al.
+2013) dataset for deblurring, the IVC-Dehazing (Ma et al.
+2015) dataset for dehazing, the IVIPC-DQA (Wu et al.
+2019) dataset for deraining, and the QADS (Zhou et al.
+2019) and SRIQA-Bench (Chen et al.
+2025) datasets for super-resolution.
+Additionally, there are datasets that encompass distortions from various algorithms, such as PIPAL (Jinjin et al.
+2020), which includes distortion types like denoising, super-resolution, and super-resolution with post-denoising, KADID-10k (Lin et al.
+2019), which contains various image distortions including denoising degradations and DiffIQA(Chen et al.
+2025), which accounts for degradation artifacts induced by diffusion-based image enhancement.
 A brief comparison of FGRestore and previous IR evaluation datasets is illustrated in Table 1.
@@
 Image Quality Assessment
-Existing image restoration (IR) algorithms typically employ full-reference image quality assessment (FR-IQA) metrics for performance evaluation, such as the conventional PSNR and SSIM (SSIM), as well as deep learning-based metrics like LPIPS (LPIPS) and DISTS (DISTS).
-To overcome the flawed ”perfect reference” assumption in traditional FR-IQA, which is challenged by imaging system limitations and superior generative methods, A-FINE (chen2025toward) proposes a generalized model that adaptively assesses both the fidelity and naturalness of test images.
+Existing image restoration (IR) algorithms typically employ full-reference image quality assessment (FR-IQA) metrics for performance evaluation, such as the conventional PSNR and SSIM (Wang et al.
+2004), as well as deep learning-based metrics like LPIPS (Zhang et al.
+2018a) and DISTS (Ding et al.
+2020).
+To overcome the flawed ”perfect reference” assumption in traditional FR-IQA, which is challenged by imaging system limitations and superior generative methods, A-FINE (Chen et al.
+2025) proposes a generalized model that adaptively assesses both the fidelity and naturalness of test images.
 No-reference IQA (NR-IQA) methods have been proposed to assess image quality without reference images.
-Early NR-IQA methods, such as NIQE(NIQE), primarily rely on hand-crafted Natural Scene Statistics (NSS) features.
+Early NR-IQA methods, such as NIQE(Mittal et al.
+2012b), primarily rely on hand-crafted Natural Scene Statistics (NSS) features.
 With the advancement of deep learning, many works have turned to neural networks for NR-IQA.
-For instance, MetaIQA(MetaIQA)leverages meta-learning to enhance model generalization.
-Additionally, methods such as CLIP-IQA(CLIP-IQA)leverage text prompts to enhance perceptual capabilities(sheng2023aesclip; yang2025language; li2025cgiaa).
-Given the powerful learning capabilities of large multi-modality models (LMMs)(sheng2025instructcrop; sheng2025agiqa), emerging work, such as Q-Align(Q-align), has begun to apply these models to quality assessment.
-Most recently, methods represented by Co-Instruct(Co-Instruct)have extended LMMs to compare multiple images.
-Furthermore, VisualQuality-R1(wu2025visualquality) introduces a novel methodology for IQA that leverages LMMs within a reinforcement learning-to-rank framework.
+For instance, MetaIQA(Zhu et al.
+2020)leverages meta-learning to enhance model generalization.
+Additionally, methods such as CLIP-IQA(Wang et al.
+2023)leverage text prompts to enhance perceptual capabilities(Sheng et al.
+2023; Yang et al.
+2025; Li et al.
+2025).
+Given the powerful learning capabilities of large multi-modality models (LMMs)(Sheng et al.
+2025a, b), emerging work, such as Q-Align(Wu et al.
+2023), has begun to apply these models to quality assessment.
+Most recently, methods represented by Co-Instruct(Wu et al.
+2024)have extended LMMs to compare multiple images.
+Furthermore, VisualQuality-R1(Wu et al.
+2025) introduces a novel methodology for IQA that leverages LMMs within a reinforcement learning-to-rank framework.
 Although these datasets and methods have contributed significantly to the IQA community, most methods are score-based and cannot effectively handle fine-grained perceptual differences brought by image restoration algorithms.
@@
 Zhang et al.
-(zhang2019fine; zhang2022fine) introduced fine-grained assessment for compressed images, focusing on distinguishing compression artifacts within similar quality levels.
+(Zhang et al.
+2018b, 2021) introduced fine-grained assessment for compressed images, focusing on distinguishing compression artifacts within similar quality levels.
 However, these prior works primarily target image compression, whereas our work specifically addresses the unique challenges of fine-grained quality assessment in image restoration tasks, where the evaluation focus shifts from compression artifacts to perceptual restoration quality differences.
@@
 Our evaluation encompasses both full-reference (FR) and no-reference (NR) methods.
-FR methods include PSNR, SSIM (SSIM), LPIPS (LPIPS), and DISTS (DISTS).
-NR methods include traditional approaches such as NIQE (NIQE), IL-NIQE (IL-NIQE), and BRISQUE (BRISQUE), as well as deep learning-based methods including DB-CNN (DBCNN), HyperIQA (HyperIQA), MetaIQA (MetaIQA), LIQE (LIQE), CLIP-IQA (CLIP-IQA), Q-Align (Q-align), and DeQA-Score (DeQA).
+FR methods include PSNR, SSIM (Wang et al.
+2004), LPIPS (Zhang et al.
+2018a), and DISTS (Ding et al.
+2020).
+NR methods include traditional approaches such as NIQE (Mittal et al.
+2012b), IL-NIQE (Zhang et al.
+2015), and BRISQUE (Mittal et al.
+2012a), as well as deep learning-based methods including DB-CNN (Zhang et al.
+2020), HyperIQA (Su et al.
+2020), MetaIQA (Zhu et al.
+2020), LIQE (Zhang et al.
+2023), CLIP-IQA (Wang et al.
+2023), Q-Align (Wu et al.
+2023), and DeQA-Score (You et al.
+2025).
 Specifically, we partition the quality score range into several intervals to examine how IQA methods perform when evaluating images with similar quality levels, which is particularly relevant for fine-grained assessment scenarios in image restoration.
-Table 2 presents the performance comparison across different MOS ranges on the PIPAL dataset (PIPAL).
+Table 2 presents the performance comparison across different MOS ranges on the PIPAL dataset (Jinjin et al.
+2020).
 While most IQA methods achieve reasonable overall SRCC values on the whole dataset, their performance dramatically deteriorates when evaluated within narrow quality ranges.
@@
 Image and Pair Collection
-We collect images from multiple restoration-specific datasets to cover different degradation and visual appearances: 1) deraining images from IVIPC-DQA (IVIPC_DQA); 2) deblurring images from MDD13 (MDD13); 3) dehazing images from IVC-Dehazing (IVC-Dehazing) and exBeDDE (exBeDDE); 4) denoising images from PIPAL (PIPAL) and KADID-10k (KADID-10k); 5) super-resolution images from QADS (QADS), RealSRQ (RealSRQ), SISRSet (SISRSet), and PIPAL (PIPAL); 6) mixture restoration images from PIPAL (PIPAL).
+We collect images from multiple restoration-specific datasets to cover different degradation and visual appearances: 1) deraining images from IVIPC-DQA (Wu et al.
+2019); 2) deblurring images from MDD13 (Liu et al.
+2013); 3) dehazing images from IVC-Dehazing (Ma et al.
+2015) and exBeDDE (Zhao et al.
+2020); 4) denoising images from PIPAL (Jinjin et al.
+2020) and KADID-10k (Lin et al.
+2019); 5) super-resolution images from QADS (Zhou et al.
+2019), RealSRQ (Jiang et al.
+2022), SISRSet (Shi et al.
+2019), and PIPAL (Jinjin et al.
+2020); 6) mixture restoration images from PIPAL (Jinjin et al.
+2020).
 After image collection, we generate image pairs by randomly pairing all images with identical content and restoration tasks, resulting in a total of 1,275,297 image pairs.
@@
 The JND represents the minimum perceptual threshold for noticeable quality differences, making it ideal for identifying imperceptible quality pairs.
-Our JND-based filtering process operates as follows: (1) randomly select 20,000 images from the collected datasets, (2) compute the JND map for each image using established JND estimation model (wu2017enhanced), and (3) generate JND noise images by adding these JND maps as noise to the original images:
+Our JND-based filtering process operates as follows: (1) randomly select 20,000 images from the collected datasets, (2) compute the JND map for each image using established JND estimation model (Wu et al.
+2017), and (3) generate JND noise images by adding these JND maps as noise to the original images:
 $I_{\text{JND}}=I+\text{JND}(I),$ (2)
 where $I$ is the original image and $\text{JND}(I)$ is the corresponding JND noise map.
-We calculate SSIM (SSIM) between each $I$ and $I_{\text{JND}}$, and use the median SSIM value $\text{SSIM}_{\text{med}}$ as our filtering threshold.
+We calculate SSIM (Wang et al.
+2004) between each $I$ and $I_{\text{JND}}$, and use the median SSIM value $\text{SSIM}_{\text{med}}$ as our filtering threshold.
 The unnoticeable pair filtering criterion is then defined as:
@@
 After applying both filtering steps, we retain 30,886 fine-grained image pairs and their corresponding 18,408 images, forming the core of our FGRestore dataset.
-Due to space limitations in the main paper, we are unable to clearly demonstrate fine-grained restoration image pair examples at sufficient resolution.
-Additional examples of fine-grained restoration image pairs are provided in the supplementary material.
+Due to space limitations, high-resolution fine-grained restoration examples are provided in the supplementary material.
+Figure 4: Consistency analysis between MOS scores and human preference rankings.
+Point sizes represent frequency of image pairs.
+Red regions indicate inconsistent cases where MOS scores and human preferences disagree.
 In-lab Subjective Study
@@
 Dataset Analysis
-To validate the necessity of fine-grained pairwise ranking annotations, we conduct a consistency analysis between MOS-based scoring and human preference ranking.
-We create a scatter plot comparing normalized MOS differences ($x$-axis) with human preference annotations ($y$-axis).
-As shown in Figure 4, the scatter plot reveals significant inconsistencies between MOS scores and human preferences.
-Points in the red regions represent inconsistent annotations where MOS scores and ranking preferences disagree.
-Critically, image pairs with smaller MOS differences exhibit larger scatter points, indicating that inconsistencies become more severe as quality differences become subtler.
-This strongly supports the necessity of fine-grained pairwise ranking, as MOS-based scoring becomes unreliable for discriminating subtle quality differences.
-Figure 4: Consistency analysis between MOS scores and human preference rankings.
-Point sizes represent frequency of image pairs.
-Red regions indicate inconsistent cases where MOS scores and human preferences disagree.
+To validate the necessity of fine-grained pairwise ranking, we analyze the consistency between MOS-based scoring and human preferences.
+Figure 4 compares normalized MOS differences ($x$-axis) with human preferences ($y$-axis), revealing significant inconsistencies.
+Points in the red regions represent cases where MOS scores and preferences disagree.
+Critically, image pairs with smaller MOS differences exhibit larger scatter, indicating that inconsistencies worsen as quality differences become subtler.
+This supports the necessity of fine-grained pairwise ranking, as MOS scoring becomes unreliable for discriminating subtle quality differences.
 Figure 5: Overview of the proposed FGResQ framework.
@@
 We compare the proposed FGResQ model against state-of-the-art FR and NR IQA methods.
-For fair comparison, all learning-based methods with public source codes are retrained using the same training-testing protocol with our scene-aware loss replacing their original regression loss, except for Compare2Score (Compare2Score) which operates in a zero-shot manner without training.
+For fair comparison, all learning-based methods with public source codes are retrained using the same training-testing protocol with our scene-aware loss replacing their original regression loss, except for Compare2Score (Zhu et al.
+2024) which operates in a zero-shot manner without training.
 We evaluate performance using SRCC and PLCC for quality score prediction, and accuracy (ACC) for fine-grained pairwise ranking.
@@
 Implementation Details
-We use CLIP ViT-B/16 as the backbone encoder for both degradation-aware feature learning and dual-branch quality prediction.
-Training images are resized to $256\times 256$ pixels and randomly cropped to $224\times 224$ patches.
+We use CLIP ViT-B/16 as the backbone encoder for degradation-aware feature learning and dual-branch quality prediction.
 The model is trained on NVIDIA RTX4090 GPUs using the Adam optimizer with cosine annealing.
 The maximum learning rate is $5\times 10^{-6}$ and the batch size is 64.
-To ensure training stability, we employ scene-aware sampling strategy that groups samples from the same source dataset within each batch.
+To ensure training stability, we employ scene-aware sampling that groups samples from the same source dataset within each batch.
 $\lambda_{1}$ and $\lambda_{2}$ are set to 5 and 1, respectively.
@@
 While some methods achieve reasonable regression performance, their fine-grained ranking capabilities remain limited.
-Figure 6: Qualitative comparison. (Best viewed zoomed in.)
 Qualitative Analysis.
@@
 Additional qualitative examples are provided in the supplementary material.
+Figure 6: Qualitative comparison. (Best viewed zoomed in.)
 Conclusion
@@
 This work is supported by National Natural Science Foundation of China under Grants 62471349, 62171340 and 62301378, Fundamental Research Funds for the Central Universities under Grant QTZX25076, and partly supported by the China Postdoctoral Science Foundation under Grant 2024M762553.
-Supplementary Material for
-Fine-grained Image Quality Assessment for Perceptual Image Restoration
-This supplementary document provides comprehensive experimental evidence, detailed analyses, and extensive visual demonstrations that substantiate the findings presented in our main paper.
-The material herein offers deeper insights into the fine-grained challenges of image restoration quality assessment and validates the effectiveness of our proposed approach across diverse scenarios.
-This document encompasses the following components:
-• Detailed Ablation Study Results: Systematic investigation of our degradation-aware feature learning (DFL) module’s contribution, demonstrating consistent improvements across multiple evaluation metrics (SRCC, PLCC, ACC) and providing empirical justification for our architectural design choices.
-• Comprehensive Performance Analysis Across Quality Ranges: Exhaustive evaluation of state-of-the-art methods across 13 datasets, revealing consistent performance degradation of existing approaches within narrow MOS score intervals and highlighting the fundamental challenges in fine-grained quality assessment.
-• Exemplars of Fine-Grained Quality Differences: Carefully curated image pairs demonstrating subtle quality variations across six restoration tasks - deblurring, denoising, deraining, super-resolution, dehazing, and mixture restoration.
-These examples illustrate the perceptual challenges that make fine-grained assessment particularly demanding for image quality assessment.
-• Thorough Qualitative Analysis: Side-by-side visual comparisons showcasing how different quality assessment methods respond to restoration results with nuanced quality differences.
-These comparisons reveal the limitations of traditional metrics and the inconsistent behavior of deep learning-based approaches in fine-grained evaluation scenarios, while demonstrating our proposed FGResQ method in accurately identifying quality preferences that align with human perceptual judgments.
-All source code and implementation details are provided in the supplementary material.
-Appendix S1 Ablation Study
-As shown in Table 3 in main paper, , unlike task-specific methods that excel in particular scenarios, FGResQ maintains robust performance across all six IR tasks, suggesting that incorporating degradation knowledge enables more generalizable quality assessment capabilities.To further validate the effectiveness of our degradation-aware feature learning (DFL) module, we conduct ablation experiments.
-Table S1 presents the average performance across all IR tasks, comparing our full FGResQ model with a variant that removes the degradation-aware learning component (w/o DFL).
-Table S1: Ablation study.
-Method SRCC PLCC ACC
-w/o DFL 0.698 0.711 0.743
-FGResQ 0.703 0.717 0.752
-Appendix S2 Comprehensive Performance Analysis Across Quality Ranges
-Following our methodology in the main paper, we partition quality score ranges into discrete intervals, and systematically evaluate existing IQA methods across 13 benchmark datasets.
-This comprehensive analysis encompasses both full-reference (FR) methods including PSNR, SSIM, LPIPS, and DISTS, as well as no-reference (NR) methods ranging from traditional approaches (NIQE, IL-NIQE, BRISQUE) to state-of-the-art learning-based methods (DB-CNN, HyperIQA, MetaIQA, LIQE, CLIP-IQA, Q-Align, DeQA-Score).
-Tables 2 through 14 present comprehensive results across two categories of datasets.
-General IQA datasets include TID2013, CSIQ, LIVE, KADID-10k and VCL@FER for synthetic distortions.
-Image restoration quality assessment datasets encompass PIPAL for super-resolution, denoise and mixture restoration, MDD13 for deblurring evaluation, exBeDDE and IVCDehazing for dehazing assessment, IVIPC-DQA for deraining quality evaluation, and super-resolution datasets QADS, RealSRQ, and SISRSet covering different scenarios.
-The results consistently demonstrate the same pattern observed in our main analysis: while most methods achieve reasonable overall correlation coefficients when evaluated on complete datasets, their performance dramatically deteriorates within narrow quality ranges.
-This reveals the inability of traditional methods to effectively conduct fine-grained image quality assessment.
-Table S2: Performance analysis across different MOS score ranges on PIPAL dataset. Methods are evaluated on subsets with MOS scores in specific ranges.
-Type Method [0.0,0.2) [0.2,0.4) [0.4,0.6) [0.6,0.8) [0.8,1.0] [0.0,1.0]
-SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC
-FR PSNR 0.323 0.266 0.082 0.107 0.209 0.224 0.161 0.135 0.072 0.021 0.422 0.420
-SSIM 0.293 0.243 0.108 0.101 0.258 0.273 0.254 0.195 0.049 0.076 0.530 0.516
-LPIPS -0.034 0.025 0.077 0.080 0.325 0.323 0.287 0.267 0.124 0.096 0.612 0.595
-DISTS 0.168 0.134 0.159 0.165 0.310 0.304 0.242 0.227 0.165 0.128 0.585 0.584
-NR NIQE -0.126 -0.101 -0.002 0.036 0.107 0.114 0.001 -0.096 0.080 0.038 0.153 0.018
-IL-NIQE -0.235 -0.144 -0.098 -0.093 0.126 0.128 0.128 0.141 0.054 0.053 0.289 0.274
-BRISQUE -0.142 -0.105 0.025 0.014 0.125 0.100 0.035 -0.048 0.131 0.096 0.185 0.078
-DB-CNN -0.157 -0.072 0.321 0.316 0.353 0.362 0.330 0.354 -0.016 0.060 0.636 0.669
-HyperIQA 0.100 0.174 0.274 0.259 0.314 0.307 0.292 0.298 0.032 -0.045 0.584 0.611
-MetaIQA 0.037 0.074 0.160 0.163 0.204 0.196 0.174 0.178 -0.101 -0.060 0.423 0.432
-LIQE -0.232 -0.035 0.053 0.059 0.175 0.084 0.299 0.240 0.107 0.182 0.479 0.318
-CLIP-IQA -0.152 -0.130 0.211 0.190 0.238 0.236 0.293 0.299 0.071 0.049 0.530 0.542
-Q-Align 0.230 0.271 0.301 0.333 0.337 0.324 0.213 0.220 0.178 0.154 0.418 0.410
-DeQA-Score 0.568 0.652 0.676 0.672 0.623 0.636 0.516 0.562 0.350 0.339 0.747 0.777
-Table S3: Performance comparison across different MOS ranges on TID2013 dataset. FR and NR represent full-reference and no-reference methods, respectively.
-Type Method [0.0,0.2) [0.2,0.4) [0.4,0.6) [0.6,0.8) [0.8,1.0] [0.0,1.0]
-SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC
-FR PSNR 0.064 -0.275 0.164 0.213 0.303 0.330 0.329 0.327 -0.033 -0.126 0.670 0.651
-SSIM -0.018 -0.197 0.192 0.262 0.194 0.178 0.380 0.371 0.175 0.113 0.596 0.633
-LPIPS -0.318 -0.010 0.308 0.371 0.269 0.252 0.407 0.356 -0.090 -0.045 0.724 0.731
-DISTS -0.136 0.052 0.345 0.419 0.268 0.268 0.343 0.350 0.026 0.060 0.688 0.736
-NR NIQE -0.191 -0.119 0.221 0.176 0.054 0.076 0.100 0.151 0.014 -0.003 0.313 0.360
-IL-NIQE -0.027 0.004 0.236 0.272 0.158 0.157 0.305 0.288 -0.141 -0.068 0.520 0.597
-BRISQUE -0.227 -0.246 0.267 0.210 0.091 0.106 0.195 0.196 0.136 0.160 0.370 0.473
-DB-CNN 0.636 0.654 0.621 0.591 0.649 0.636 0.496 0.507 0.293 0.303 0.889 0.905
-HyperIQA 0.391 0.453 0.587 0.605 0.391 0.347 0.339 0.333 0.000 0.027 0.810 0.839
-MetaIQA 0.400 0.502 0.595 0.618 0.563 0.546 0.560 0.526 0.529 0.631 0.885 0.875
-LIQE -0.018 -0.140 0.484 0.321 0.468 0.335 0.513 0.508 0.322 0.375 0.869 0.797
-CLIP-IQA 0.582 0.567 0.438 0.446 0.467 0.495 0.403 0.413 0.087 0.040 0.813 0.835
-Q-Align 0.346 0.464 0.489 0.497 0.510 0.451 0.469 0.472 0.303 0.349 0.867 0.870
-DeQA-Score 0.382 0.448 0.566 0.560 0.572 0.566 0.514 0.525 0.467 0.429 0.891 0.898
-Table S4: Performance comparison across different MOS ranges on CSIQ dataset. FR and NR represent full-reference and no-reference methods, respectively.
-Type Method [0.0,0.2) [0.2,0.4) [0.4,0.6) [0.6,0.8) [0.8,1.0] [0.0,1.0]
-SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC
-FR PSNR 0.006 0.236 0.536 0.564 0.010 -0.003 0.305 0.411 0.534 0.620 0.808 0.801
-SSIM 0.103 -0.035 0.409 0.455 -0.040 0.012 0.378 0.355 0.646 0.475 0.850 0.793
-LPIPS 0.333 0.250 0.441 0.461 0.343 0.331 0.620 0.582 0.657 0.568 0.925 0.903
-DISTS 0.055 0.037 0.664 0.661 0.288 0.334 0.508 0.509 0.673 0.593 0.933 0.929
-NR NIQE 0.661 0.501 0.424 0.351 -0.320 -0.161 0.316 0.346 0.296 0.260 0.619 0.700
-IL-NIQE 0.346 0.086 0.446 0.395 0.367 0.318 0.524 0.428 0.216 0.193 0.809 0.787
-BRISQUE 0.200 0.215 0.385 0.278 0.076 0.064 0.302 0.301 0.146 0.172 0.559 0.688
-DB-CNN 0.624 0.605 0.696 0.657 0.693 0.638 0.526 0.550 0.448 0.483 0.929 0.942
-HyperIQA 0.730 0.629 0.472 0.408 0.340 0.419 0.566 0.575 0.316 0.315 0.858 0.891
-MetaIQA 0.640 0.605 0.624 0.575 0.552 0.493 0.477 0.417 0.469 0.450 0.888 0.882
-LIQE 0.566 0.476 0.579 0.433 0.469 0.489 0.686 0.687 0.675 0.664 0.947 0.888
-CLIP-IQA 0.697 0.591 0.306 0.372 0.165 0.203 0.488 0.498 0.352 0.367 0.842 0.868
-Q-Align 0.649 0.549 0.576 0.560 0.409 0.444 0.637 0.607 0.341 0.370 0.877 0.904
-DeQA-Score 0.830 0.586 0.521 0.491 0.528 0.569 0.546 0.560 0.552 0.551 0.938 0.952
-Table S5: Performance comparison across different MOS ranges on LIVE dataset. FR and NR represent full-reference and no-reference methods, respectively.
-Type Method [0.0,0.2) [0.2,0.4) [0.4,0.6) [0.6,0.8) [0.8,1.0] [0.0,1.0]
-SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC
-FR PSNR 0.687 0.627 0.296 0.233 0.525 0.494 0.618 0.628 0.882 0.922 0.952 0.806
-SSIM 0.685 0.654 0.298 0.326 0.624 0.637 0.706 0.682 0.894 0.858 0.971 0.885
-LPIPS 0.710 0.655 0.263 0.225 0.625 0.641 0.669 0.639 0.891 0.847 0.960 0.845
-DISTS 0.710 0.635 0.286 0.215 0.728 0.702 0.703 0.719 0.880 0.879 0.979 0.936
-NR NIQE 0.591 0.610 0.151 0.100 0.442 0.449 0.677 0.581 0.474 0.535 0.931 0.645
-IL-NIQE 0.677 0.612 0.220 0.113 0.610 0.619 0.501 0.540 0.297 0.196 0.921 0.690
-BRISQUE 0.576 0.546 0.506 0.543 0.765 0.768 0.864 0.808 0.541 0.515 0.971 0.976
-DB-CNN 0.476 0.486 0.486 0.615 0.797 0.770 0.816 0.780 0.666 0.676 0.977 0.982
-HyperIQA 0.676 0.694 0.605 0.715 0.813 0.803 0.856 0.834 0.625 0.655 0.977 0.986
-MetaIQA 0.786 0.604 0.581 0.589 0.494 0.462 0.284 0.157 0.266 0.131 0.874 0.837
-LIQE -0.185 -0.169 0.461 0.397 0.791 0.769 0.693 0.691 0.365 0.287 0.959 0.943
-CLIP-IQA 0.428 0.410 0.452 0.518 0.629 0.615 0.736 0.700 0.500 0.541 0.961 0.968
-Q-Align 0.889 0.994 0.907 0.953 0.803 0.716 0.891 0.926 0.554 0.972 0.923 0.928
-DeQA-Score 0.845 0.987 0.895 0.922 0.773 0.842 0.895 0.933 0.877 0.991 0.974 0.976
-Table S6: Performance analysis across different MOS score ranges on KADID-10K dataset. Methods are evaluated on subsets with MOS scores in specific ranges.
-Type Method [0.0,0.2) [0.2,0.4) [0.4,0.6) [0.6,0.8) [0.8,1.0] [0.0,1.0]
-SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC
-FR PSNR 0.409 0.370 0.212 0.219 0.147 0.126 0.163 0.168 0.162 0.153 0.675 0.554
-SSIM 0.442 0.371 0.210 0.176 0.137 0.176 0.180 0.152 0.288 0.293 0.609 0.581
-LPIPS 0.416 0.323 0.355 0.298 0.224 0.231 0.276 0.233 0.412 0.399 0.825 0.752
-DISTS 0.365 0.315 0.312 0.285 0.244 0.241 0.311 0.301 0.473 0.458 0.807 0.799
-NR NIQE 0.036 0.043 0.091 0.076 0.043 0.058 0.061 0.051 0.366 0.337 0.386 0.410
-IL-NIQE -0.059 -0.048 0.095 0.111 0.057 0.089 0.160 0.192 0.290 0.302 0.526 0.543
-BRISQUE 0.082 0.039 0.045 0.020 0.048 0.037 0.041 0.047 0.256 0.272 0.323 0.387
-DB-CNN 0.004 0.009 0.255 0.248 0.311 0.298 0.382 0.386 0.337 0.327 0.822 0.809
-HyperIQA 0.047 0.022 0.275 0.256 0.310 0.306 0.373 0.362 0.519 0.460 0.803 0.806
-MetaIQA 0.023 0.006 0.159 0.178 0.228 0.208 0.367 0.362 0.487 0.474 0.793 0.782
-LIQE 0.194 0.145 0.318 0.284 0.343 0.335 0.359 0.270 0.285 0.195 0.798 0.758
-CLIP-IQA 0.125 0.110 0.228 0.235 0.217 0.215 0.313 0.317 0.374 0.346 0.790 0.790
-Q-Align 0.436 0.426 0.488 0.482 0.534 0.502 0.534 0.436 0.604 0.445 0.958 0.962
-DeQA-Score 0.468 0.532 0.638 0.622 0.673 0.742 0.689 0.762 0.726 0.753 0.926 0.923
-Table S7: Performance analysis across different MOS score ranges on VCL@FER. Methods are evaluated on subsets with MOS scores in specific ranges.
-Type Method [0.0,0.2) [0.2,0.4) [0.4,0.6) [0.6,0.8) [0.8,1.0] [0.0,1.0]
-SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC
-FR PSNR 0.114 0.111 0.159 0.178 0.316 0.287 0.422 0.429 0.413 0.488 0.846 0.857
-SSIM 0.414 0.480 0.483 0.484 0.398 0.432 0.488 0.381 0.063 0.242 0.919 0.812
-LPIPS 0.569 0.606 0.575 0.569 0.449 0.423 0.488 0.358 0.301 0.289 0.896 0.808
-DISTS 0.546 0.650 0.630 0.650 0.532 0.562 0.545 0.530 0.462 0.329 0.942 0.913
-NR NIQE 0.185 0.089 0.461 0.414 0.413 0.373 0.515 0.439 -0.119 0.174 0.835 0.549
-IL-NIQE 0.072 0.100 0.516 0.414 0.306 0.257 0.403 0.385 0.301 0.110 0.796 0.761
-BRISQUE 0.274 0.325 0.525 0.409 0.607 0.600 0.448 0.451 0.601 0.571 0.917 0.904
-DB-CNN 0.654 0.628 0.733 0.689 0.577 0.627 0.695 0.713 0.273 0.239 0.960 0.959
-HyperIQA 0.725 0.711 0.779 0.740 0.558 0.608 0.647 0.612 -0.112 -0.134 0.946 0.945
-MetaIQA 0.336 0.384 0.277 0.293 0.233 0.110 0.216 0.169 0.123 -0.115 0.687 0.635
-LIQE 0.728 0.400 0.745 0.742 0.630 0.677 0.654 0.612 0.406 0.493 0.963 0.954
-CLIP-IQA 0.345 0.291 0.547 0.542 0.356 0.372 0.708 0.696 0.371 0.389 0.952 0.958
-Q-Align 0.718 0.275 0.576 0.629 0.435 0.380 0.633 0.692 0.594 0.607 0.890 0.858
-DeQA-Score 0.631 0.631 0.720 0.782 0.678 0.717 0.812 0.809 0.559 0.661 0.957 0.956
-Table S8: Performance analysis across different MOS score ranges on exBeDDE. Methods are evaluated on subsets with MOS scores in specific ranges.
-Type Method [0.0,0.2) [0.2,0.4) [0.4,0.6) [0.6,0.8) [0.8,1.0] [0.0,1.0]
-SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC
-FR PSNR -0.235 -0.234 0.077 0.049 -0.290 -0.225 -0.304 -0.322 -0.233 -0.309 -0.705 -0.701
-SSIM -0.176 -0.177 0.002 0.002 -0.027 0.017 -0.311 -0.320 -0.455 -0.491 -0.730 -0.720
-LPIPS -0.075 -0.055 -0.117 -0.189 -0.421 -0.450 -0.245 -0.246 -0.321 -0.403 -0.747 -0.752
-DISTS 0.001 -0.060 -0.190 -0.236 -0.322 -0.285 -0.404 -0.341 -0.263 -0.227 -0.783 -0.755
-NR NIQE 0.123 0.091 -0.082 -0.065 0.198 0.133 -0.128 -0.145 -0.050 -0.055 0.029 0.031
-IL-NIQE 0.108 0.082 -0.036 -0.039 0.032 0.061 0.014 -0.032 -0.402 -0.335 0.107 0.074
-BRISQUE -0.086 -0.091 0.116 0.140 0.180 0.195 -0.223 -0.237 0.241 0.220 0.118 0.068
-DB-CNN 0.155 0.197 0.344 0.371 0.200 0.264 0.396 0.384 0.731 0.721 0.899 0.905
-HyperIQA 0.125 0.191 0.324 0.298 0.226 0.206 0.159 0.143 0.594 0.610 0.847 0.863
-MetaIQA 0.135 0.239 0.297 0.298 0.262 0.263 0.223 0.221 0.662 0.617 0.860 0.881
-LIQE 0.017 0.143 0.433 0.407 0.383 0.356 0.385 0.376 0.734 0.492 0.895 0.893
-CLIP-IQA 0.168 0.179 0.397 0.406 0.328 0.342 0.215 0.217 0.524 0.491 0.879 0.903
-Q-Align 0.013 0.044 0.153 0.170 0.218 0.260 0.286 0.289 0.472 0.432 0.814 0.815
-DeQA-Score 0.105 0.142 0.201 0.186 0.172 0.275 0.258 0.265 0.707 0.727 0.863 0.879
-Table S9: Performance analysis across different MOS score ranges on IVCDehazing. Methods are evaluated on subsets with MOS scores in specific ranges.
-Type Method [0.0,0.2) [0.2,0.4) [0.4,0.6) [0.6,0.8) [0.8,1.0] [0.0,1.0]
-SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC
-FR PSNR - - - - - - - - - - - -
-SSIM - - - - - - - - - - - -
-LPIPS - - - - - - - - - - - -
-DISTS - - - - - - - - - - - -
-NR NIQE -1.000 -1.000 -0.113 -0.056 0.127 0.059 0.193 0.150 0.600 0.920 0.357 0.455
-IL-NIQE 1.000 1.000 -0.042 -0.162 0.582 0.440 0.132 0.105 -0.200 -0.576 0.381 0.509
-BRISQUE 1.000 1.000 0.459 0.407 -0.291 -0.022 0.062 -0.003 0.200 0.582 0.122 0.124
-DB-CNN 1.000 1.000 -0.127 -0.175 0.255 0.268 0.324 0.397 -0.200 0.310 0.753 0.654
-HyperIQA 1.000 1.000 -0.177 -0.101 0.800 0.715 0.130 0.002 -0.400 -0.667 0.404 0.592
-MetaIQA 1.000 1.000 0.127 0.150 -0.346 -0.422 0.171 0.270 0.000 0.085 0.683 0.618
-LIQE 1.000 1.000 -0.438 -0.513 0.255 0.455 0.443 0.380 0.200 0.377 0.630 0.540
-CLIP-IQA 1.000 1.000 0.223 0.132 -0.255 -0.332 0.107 0.179 0.000 0.154 0.530 0.559
-Q-Align 1.000 1.000 0.329 0.278 0.143 0.072 0.600 0.634 -1.000 -0.978 0.328 0.303
-DeQA-Score 1.000 1.000 0.410 0.532 0.643 0.778 -0.150 -0.323 -0.400 -0.460 0.103 0.177
-Table S10: Performance analysis across different MOS score ranges on IVIPC-DQA dataset. Methods are evaluated on subsets with MOS scores in specific ranges.
-Type Method [0.0,0.2) [0.2,0.4) [0.4,0.6) [0.6,0.8) [0.8,1.0] [0.0,1.0]
-SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC
-FR PSNR - - - - - - - - - - - -
-SSIM - - - - - - - - - - - -
-LPIPS - - - - - - - - - - - -
-DISTS - - - - - - - - - - - -
-NR NIQE 0.255 0.240 0.077 0.073 0.121 0.077 - - - - 0.032 0.008
-IL-NIQE -0.094 -0.073 0.038 0.045 0.184 0.126 - - - - 0.002 0.033
-BRISQUE -0.003 -0.115 -0.018 -0.028 0.113 0.172 - - - - 0.090 0.094
-DB-CNN 0.164 0.165 0.338 0.369 0.238 0.264 - - - - 0.491 0.525
-HyperIQA 0.011 0.038 0.369 0.417 0.050 0.111 - - - - 0.475 0.503
-MetaIQA -0.022 -0.040 0.393 0.353 0.319 0.364 - - - - 0.550 0.546
-LIQE -0.098 0.044 0.329 0.332 0.136 0.152 - - - - 0.483 0.479
-CLIP-IQA -0.043 -0.041 0.159 0.189 0.215 0.240 - - - - 0.400 0.450
-Q-Align 0.015 -0.105 0.272 0.349 0.245 0.324 - - - - 0.255 0.302
-DeQA-Score 0.571 0.544 0.512 0.602 0.506 0.502 - - - - 0.533 0.581
-Table S11: Performance analysis across different MOS score ranges on MDD13 dataset. Methods are evaluated on subsets with MOS scores in specific ranges.
-Type Method [0.0,0.2) [0.2,0.4) [0.4,0.6) [0.6,0.8) [0.8,1.0] [0.0,1.0]
-SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC
-FR PSNR 1.000 0.656 0.182 0.154 0.321 0.317 0.048 0.066 0.014 -0.068 0.186 0.168
-SSIM 1.000 0.657 0.015 -0.008 0.261 0.294 0.289 0.167 0.125 0.071 0.445 0.353
-LPIPS 0.800 0.799 0.232 0.093 0.252 0.091 0.340 0.326 0.472 0.451 0.777 0.702
-DISTS 1.000 0.922 0.400 0.336 0.295 0.292 0.600 0.591 0.790 0.820 0.907 0.901
-NR NIQE -0.200 -0.702 0.296 0.302 0.128 0.097 0.253 0.287 0.372 0.475 0.386 0.413
-IL-NIQE 0.000 -0.287 0.331 0.127 0.176 0.150 0.258 0.287 0.394 0.382 0.591 0.515
-BRISQUE 0.200 0.029 -0.130 -0.238 0.091 0.075 0.304 0.304 0.206 0.227 0.356 0.356
-DB-CNN 0.800 0.698 0.540 0.671 0.437 0.447 0.538 0.551 0.476 0.515 0.868 0.888
-HyperIQA 0.000 0.318 0.577 0.603 0.370 0.388 0.551 0.551 0.377 0.387 0.862 0.879
-MetaIQA 1.000 0.843 0.455 0.445 0.310 0.279 0.462 0.391 0.282 0.218 0.806 0.826
-LIQE 0.949 0.656 0.536 0.559 0.497 0.442 0.594 0.582 0.370 0.400 0.899 0.814
-CLIP-IQA 0.316 0.649 0.449 0.610 0.499 0.524 0.604 0.628 0.556 0.622 0.908 0.910
-Q-Align -0.400 -0.086 0.380 0.488 0.296 0.301 0.589 0.588 0.672 0.663 0.884 0.880
-DeQA-Score 0.800 0.916 0.432 0.608 0.429 0.445 0.529 0.532 0.493 0.539 0.887 0.905
-Table S12: Performance analysis across different MOS score ranges on QADS dataset. Methods are evaluated on subsets with MOS scores in specific ranges.
-Type Method [0.0,0.2) [0.2,0.4) [0.4,0.6) [0.6,0.8) [0.8,1.0] [0.0,1.0]
-SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC
-FR PSNR 1.000 0.656 0.182 0.154 0.321 0.317 0.048 0.066 0.014 -0.068 0.186 0.168
-SSIM 1.000 0.657 0.015 -0.008 0.261 0.294 0.289 0.167 0.125 0.071 0.445 0.353
-LPIPS 0.800 0.799 0.232 0.093 0.252 0.091 0.340 0.326 0.472 0.451 0.777 0.702
-DISTS 1.000 0.922 0.400 0.336 0.295 0.292 0.600 0.591 0.790 0.820 0.907 0.901
-NR NIQE -0.200 -0.702 0.296 0.302 0.128 0.097 0.253 0.287 0.372 0.475 0.386 0.413
-IL-NIQE 0.000 -0.287 0.331 0.127 0.176 0.150 0.258 0.287 0.394 0.382 0.591 0.515
-BRISQUE 0.200 0.029 -0.130 -0.238 0.091 0.075 0.304 0.304 0.206 0.227 0.356 0.356
-DB-CNN 0.800 0.698 0.540 0.671 0.437 0.447 0.538 0.551 0.476 0.515 0.868 0.888
-HyperIQA 0.000 0.318 0.577 0.603 0.370 0.388 0.551 0.551 0.377 0.387 0.862 0.879
-MetaIQA 1.000 0.843 0.455 0.445 0.310 0.279 0.462 0.391 0.282 0.218 0.806 0.826
-LIQE 0.949 0.656 0.536 0.559 0.497 0.442 0.594 0.582 0.370 0.400 0.899 0.814
-CLIP-IQA 0.316 0.649 0.449 0.610 0.499 0.524 0.604 0.628 0.556 0.622 0.908 0.910
-Q-Align -0.400 -0.086 0.380 0.488 0.296 0.301 0.589 0.588 0.672 0.663 0.884 0.880
-DeQA-Score 0.800 0.916 0.432 0.608 0.429 0.445 0.529 0.532 0.493 0.539 0.887 0.905
-Table S13: Performance analysis across different MOS score ranges on RealSRQ dataset. Methods are evaluated on subsets with MOS scores in specific ranges.
-Type Method [0.0,0.2) [0.2,0.4) [0.4,0.6) [0.6,0.8) [0.8,1.0] [0.0,1.0]
-SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC
-FR PSNR 1.000 1.000 -0.149 -0.151 -0.069 0.010 - - - - -0.035 0.013
-SSIM 1.000 1.000 -0.173 -0.116 0.021 0.118 - - - - -0.029 0.050
-LPIPS 1.000 1.000 -0.148 -0.106 0.271 0.202 - - - - -0.003 0.071
-DISTS 1.000 1.000 -0.147 -0.132 0.311 0.309 - - - - -0.034 0.057
-NR NIQE -1.000 -1.000 -0.017 -0.005 0.359 0.362 - - - - 0.013 0.050
-IL-NIQE -1.000 -1.000 -0.031 -0.014 0.218 0.205 - - - - 0.006 0.032
-BRISQUE -1.000 -1.000 0.059 0.117 0.095 0.156 - - - - 0.045 -0.018
-DB-CNN -1.000 -1.000 0.610 0.630 0.786 0.801 - - - - 0.739 0.838
-HyperIQA -1.000 -1.000 0.474 0.490 0.705 0.743 - - - - 0.602 0.609
-MetaIQA 1.000 1.000 0.106 -0.003 -0.061 0.029 - - - - 0.171 0.134
-LIQE -1.000 -1.000 0.412 0.347 0.561 0.421 - - - - 0.537 0.525
-CLIP-IQA -1.000 -1.000 0.395 0.439 0.636 0.660 - - - - 0.555 0.659
-Q-Align -1.000 -1.000 -0.058 0.005 0.201 0.303 - - - - 0.100 0.227
-DeQA-Score -1.000 -1.000 0.267 0.269 0.685 0.679 - - - - 0.445 0.609
-Table S14: Performance analysis across different MOS score ranges on SISRSet dataset. Methods are evaluated on subsets with MOS scores in specific ranges.
-Type Method [0.0,0.2) [0.2,0.4) [0.4,0.6) [0.6,0.8) [0.8,1.0] [0.0,1.0]
-SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC
-FR PSNR -0.314 -0.168 -0.136 0.077 0.261 0.077 0.762 0.741 -0.367 -0.231 0.589 0.613
-SSIM -0.074 -0.072 -0.009 -0.068 0.029 -0.240 0.833 0.648 -0.346 -0.196 0.371 0.406
-LPIPS 0.279 0.165 -0.600 -0.511 -0.540 -0.592 -0.167 0.129 -0.217 -0.364 0.517 0.560
-DISTS 0.220 0.122 -0.409 -0.427 -0.503 -0.576 -0.048 0.242 -0.045 -0.415 0.665 0.653
-NR NIQE 0.288 0.260 0.446 0.525 0.753 0.809 -0.191 -0.238 0.351 0.430 0.680 0.683
-IL-NIQE 0.712 0.687 0.236 0.371 0.562 0.594 -0.071 -0.120 0.338 0.449 0.639 0.616
-BRISQUE 0.664 0.680 0.055 -0.064 0.213 0.444 -0.310 -0.194 0.512 0.458 0.888 0.876
-DB-CNN 0.664 0.655 0.791 0.675 0.852 0.804 0.476 0.040 0.772 0.683 0.915 0.902
-HyperIQA 0.801 0.810 0.636 0.642 0.617 0.623 0.262 0.053 0.452 0.491 0.959 0.946
-MetaIQA 0.444 0.385 0.091 0.180 0.050 0.435 -0.357 -0.477 0.479 0.222 0.638 0.511
-LIQE 0.191 0.266 0.364 0.100 0.631 0.596 -0.333 -0.476 0.397 0.122 0.307 0.436
-CLIP-IQA 0.606 0.653 0.164 0.225 0.125 0.043 0.071 0.032 0.436 0.371 0.945 0.941
-Q-Align 0.805 0.799 0.564 0.528 0.208 0.197 -0.119 -0.175 -0.204 -0.081 0.679 0.665
-DeQA-Score 0.553 0.528 0.427 0.360 0.356 0.336 -0.381 -0.282 0.591 0.191 0.496 0.417
-Appendix S3 Fine-grained Image Pairs Visualization
-This section presents visual examples demonstrating the perceptual challenges in fine-grained quality assessment across six restoration tasks.
-These examples represent the subtle quality variations commonly encountered in practical image restoration applications, where algorithmic comparisons and parameter optimization require sensitive discrimination between restoration results with marginal quality differences.
-Figures S1 through S6 showcase image pairs with subtle quality differences in deblurring, denoising, deraining, super-resolution, dehazing, and mixture restoration tasks.
-The image pairs demonstrate cases where slight variations in restoration effectiveness can significantly impact perceptual quality despite producing similar overall visual appearance.
-These carefully selected examples illustrate the inherent complexity of fine-grained quality assessment in image restoration.
-The visual evidence demonstrates that distinguishing subtle quality differences requires sophisticated understanding of restoration-specific artifacts and perceptual quality factors.
-These examples highlight the necessity for FGResQ designed to handle the nuanced quality differences encountered in image restoration applications, where traditional evaluation paradigms may not capture the full spectrum of perceptual quality variations.
-Figure S1: Fine-grained image pairs for Deblurring
-Figure S2: Fine-grained image pairs for Denoising
-Figure S3: Fine-grained image pairs for Deraining
-Figure S4: Fine-grained image pairs for Super-Resolution
-Figure S5: Fine-grained image pairs for Dehazing
-Figure S6: Fine-grained image pairs for Mixture Restoration
-Appendix S4 Qualitative Analysis
-This section provides visual evidence of how different quality assessment methods respond to fine-grained quality differences in restored images.
-Figures S7 through S12 present qualitative comparisons on representative fine-grained image pairs across different restoration tasks.
-The results demonstrate systematic limitations of existing IQA methods in fine-grained scenarios.
-Traditional metrics like PSNR and SSIM often produce nearly identical scores for images with subtle quality differences, failing to provide meaningful quality discrimination.
-Advanced learning-based methods such as CLIP-IQA and DeQA-Score show improved sensitivity compared to traditional metrics but still exhibit inconsistent behavior in fine-grained evaluation.
-These methods frequently produce incorrect rankings when quality differences become subtle, particularly in challenging scenarios like mixture restoration where multiple degradation types interact.
-In contrast, our proposed FGResQ demonstrates superior performance across all tested scenarios.
-The method consistently identifies the superior image in fine-grained pairs where existing approaches fail, particularly excelling in challenging mixture restoration and dehazing scenarios.
-FGResQ’s degradation-aware feature learning enables more nuanced understanding of restoration quality, while the dual-branch architecture effectively captures both coarse-grained and fine-grained quality relationships.
-These visual comparisons complement our quantitative analysis and provide intuitive evidence demonstrating FGResQ’s effectiveness in addressing the fundamental limitations of existing approaches for fine-grained quality assessment in image restoration applications.
-Figure S7: Qualitative comparison for Deblurring
-Figure S8: Qualitative comparison for Denoising
-Figure S9: Qualitative comparison for Deraining
-Figure S10: Qualitative comparison for Super-Resolution
-Figure S11: Qualitative comparison for Dehazing
-Figure S12: Qualitative comparison for Mixture Restoration